### v2

# Linear-Size $\mathrm{QAC}^{0}$ Channels: Learning, Testing and Hardness

Abstract
Shallow quantum circuits have attracted increasing attention in recent years, due to the fact that current noisy quantum hardware can only perform faithful quantum computation for a short amount of time. The constant-depth quantum circuits $\mathbf{QAC}^{0}$, a quantum counterpart of $\mathbf{AC}^{0}$ circuits, are the polynomial-size and constant-depth quantum circuits composed of only single-qubit unitaries and polynomial-size generalized Toffoli gates. The computational power of $\mathbf{QAC}^{0}$ has been extensively investigated in recent years [10.5555/2011679.2011682, DBLP:journals/corr/abs-2005-12169, rosenthal:LIPIcs.ITCS.2021.32, NPVY24, ADOY24].
In this paper, we are concerned with $\mathbf{QLC}^{0}$ circuits, which are linear-size $\mathbf{QAC}^{0}$ circuits, a quantum counterpart of $\mathbf{LC}^{0}$. We provide a comprehensive study of $\mathbf{QLC}^{0}$ circuits. Our results are as follows.
  • We show that depth-$d$ $\mathbf{QAC}^{0}$ circuits working on $n$ input qubits and $a$ ancilla qubits have approximate degree at most $\tilde{O}\left((n+a)^{1-2^{-d}}\right)$, improving the $\tilde{O}\left((n+a)^{1-3^{-d}}\right)$ degree upper bound of [ADOY24]. Consequently, this directly implies that to compute the parity function, $\mathbf{QAC}^{0}$ circuits need at least $\tilde{O}\left(n^{1+2^{-d}}\right)$ circuit size. We obtain this bound by improving the techniques in [ADOY24] by using the techniques of unitary dilation and operator dilation.
  • We present the first agnostic learning algorithm for $\mathbf{QLC}^{0}$ channels using subexponential running time and queries. Moreover, we also establish exponential lower bounds on the query complexity of learning $\mathbf{QAC}^{0}$ channels under both the spectral norm distance of the Choi matrix and the diamond norm distance.
  • We present a tolerant testing algorithm which determines whether an unknown quantum channel is a $\mathbf{QLC}^{0}$ channel. This tolerant testing algorithm is based on our agnostic learning algorithm.
Our approach leverages low-degree approximations of $\mathbf{QAC}^{0}$ circuits and Pauli analysis as key technical tools. Collectively, these results advance our understanding of agnostic learning for shallow quantum circuits.
1 Introduction
Current quantum hardware suffers from decoherence and is only able to perform faithful quantum computation for a short amount of time. This motivates the study of shallow quantum circuits, which have received increasing attention in recent years [bravyi2024classical, HLB+24, vasconcelos2024learning, arunachalam_et_al:LIPIcs.ICALP.2024.13]. A simple yet important class of shallow quantum circuits are the $\mathbf{QNC}^{0}$ circuits, which are polynomial size, constant depth quantum circuits with single-qubit and two-qubit quantum gates. Despite these constrains, $\mathbf{QNC}^{0}$ circuits already demonstrate computational advantages over classical algorithms in specific tasks [doi:10.1126/science.aar3106, 10.1145/3313276.3316404, bravyi2020quantum, watts2023unconditional], such as sampling from classically intractable distributions [gao2017quantum, haferkamp2020closing, coble2022quasi]. However, due to the light-cone constraint, the ability of $\mathbf{QNC}^{0}$ circuits to generate long-range entanglement or to compute Boolean functions is significantly limited.
Green, Homer, Moore, and Pollett [moore1999quantum, moor] introduced $\mathbf{QAC}^{0}$ circuits as a quantum counterpart of classical $\mathbf{AC}^{0}$ circuits, which are constant-depth polynomial-size quantum circuits consisting of single-qubit unitary gates together with arbitrarily size generalized Toffoli gates. Classical $\mathbf{AC}^{0}$ is a central object in circuit complexity, whose computational power is well understood and which delineates the frontier of current lower bound techniques. Thus, it is intriguing to investigate the computational power of $\mathbf{QAC}^{0}$. Recent progress on quantum hardware has also enabled the implementation of long-range multi-qubit operations, including the generalized Toffoli gate [rasmussen2020single, goel2021native, nikolaeva2025scalable] and the quantum fan-out gate [Gokhale2020QuantumFC], underscoring the relevance of $\mathbf{QAC}^{0}$ as a model for near-term quantum computation.
Using the Pauli analysis framework, Nadimpalli, Parham, Vasconcelos, and Yuen [NPVY24] characterized the structure of $\mathbf{QAC}^{0}$ circuits of depth $d$ and ancilla size $n^{O(1/d)}$. They showed that $\mathbf{QAC}^{0}$ circuits exhibit a low degree Pauli concentration similar to $\mathbf{AC}^{0}$ circuits when the ancilla is of size $n^{O(1/d)}$. Later, Anshu, Dong, Ou, and Yao [ADOY24] demonstrated that $\mathbf{QAC}^{0}$ circuits with barely linear-size $n^{1+3^{-d}}$ have a small approximation degree. Thus, linear-size $\mathbf{QAC}^{0}$ circuits stand at the frontier of the lower bounds of quantum circuits. This motivates us to investigate the family of quantum circuits $\mathbf{QLC}^{0}$  as a quantum analog of $\mathbf{LC}^{0}$ [1663737], which are linear size $\mathbf{QAC}^{0}$ circuits.
Hardness
A central challenge in studying $\mathbf{QAC}^{0}$ circuits is to understand the computational power of $\mathbf{QAC}^{0}$. In particular, given that $\mathbf{AC}^{0}$ circuits can not compute the parity function [10.1145/12130.12132], a core problem is whether $\mathbf{QAC}^{0}$ circuits are stronger than $\mathbf{AC}^{0}$ circuits and can compute the parity function. This problem has been extensively studied in the past [10.5555/2011679.2011682, DBLP:journals/corr/abs-2005-12169, rosenthal:LIPIcs.ITCS.2021.32, NPVY24, ADOY24], where they gave different ancilla lower bounds for $\mathbf{QAC}^{0}$ circuits to compute the parity function. Notably, given that the parity function has linear Fourier degree, the low-degree approximation results of [NPVY24] and [ADOY24] show that $\mathbf{QAC}^{0}$ circuits can not compute the parity function, if there is only $O\left(n^{1/d}\right)$ and $O\left(n^{1+3^{-d}}\right)$ ancilla, respectively. Given that $\mathbf{QAC}^{0}$ circuits with exponential ancilla can compute the parity function [rosenthal:LIPIcs.ITCS.2021.32], a still open question is what is the minimal number of ancilla qubits required to compute the parity function.
An observation of [ADOY24] shows that any ancilla lower bound of the form $n^{1+\exp\left(-o(d)\right)}$ can be boosted to $n^{c}$ for any constant $c>1$. Thus if we can slightly improve the ancilla lower bound of [ADOY24], then we can prove that any $\mathbf{QAC}^{0}$ circuit can not compute the parity function. So one possible path to resolve this open question is to improve the exponent in the $O\left(n^{1+3^{-d}}\right)$ lower bound.
PAC Learning and Agnostic Learning
The task of learning an unknown quantum process, a.k.a. Quantum process tomography, is a fundamental task in quantum physics and quantum computation [Chuang_1997, Poyatos_1997, PhysRevLett.86.4195, O_Brien_2004, Scott_2008, PhysRevA.77.032322, Haah_2023]. This problem has been extensively studied since the early days of quantum computation and has found broad applications in diverse areas, including quantum cryptography [Arapinis2021quantumphysical, doi:10.1126/science.adv8590], quantum metrology [PhysRevLett.96.010401, Giovannetti2011], and error mitigation [RevModPhys.95.045005].
Haah, Kothari, O’Donnell, and Tang [Haah_2023] have shown that learning an arbitrary quantum channel requires exponential resources. However, in practice we are not always faced with learning completely arbitrary quantum channels. Recently a number of learning algorithms for shallow quantum circuits have emerged in the literature [HLB+24, vasconcelos2024learning, arunachalam_et_al:LIPIcs.ICALP.2024.13, bao2024learning], alongside many learning algorithms for quantum channels with other specific underlying structures, such as junta channels [doi:10.1137/1.9781611977554.ch43, pmlr-v195-bao23b], and Pauli channels [Harper_2020, Flammia_2020, Harper_2021, Chen_2022, Kunjummen_2023, wadhwa2024agnostic]. In the standard model of PAC learning, it assumes that the target channels exactly satisfy the assumed structure. In real-world quantum systems, this assumption is often unrealistic due to unavoidable noise and imperfect gate operations. This motivates the study of agnostic learning, where the learning algorithm does not assume the target lies exactly within the hypothesis class.
Agnostic PAC learning, introduced by Kearns, Schapire, and Sellie [10.1145/130385.130424], is a more general and robust learning model. It seeks to find a hypothesis within a given class that best approximates an arbitrary target function. Agnostic learning has received significant attention since it was introduced. Despite its flexibility, designing agnostic learning algorithms with nontrivial performance guarantees remains a significant challenge, even when allowing sub-exponential running time [10.1145/3444815.3444825].
The quantum version of agnostic learning for functions was first investigated by Arunachalam and de Wolf [10.5555/3291125.3309633], who established tight lower bounds on the sample complexity of agnostic learning in terms of the VC dimension. More recently, agnostic learning has been extended to quantum states, including product states [10.1145/3717823.3718207], and stabilizer states [chen2024stabilizerbootstrappingrecipeefficient, grewal2025agnostictomographystabilizerproduct]. Wadhwa, Lewis, Kashefi, and Doosti [wadhwa2024agnostic] further introduced agnostic process tomography, proposing algorithms for learning several classes of quantum channels, including bounded-gate circuits, Clifford circuits, Pauli strings, $k$-juntas, and low-degree circuits.
Despite this growing body of work, relatively little is known about agnostic learning for shallow quantum circuits. To the best of our knowledge, the only existing result prior to ours in this direction is due to Wadhawan, Lewis, Kashefi, and Doosti [wadhwa2024agnostic] who initiated the study of agnostic learning in this setting. We summarize all currently known results on learning shallow quantum circuits with many-qubit gates, including both upper and lower bounds on their agnostic and non-agnostic learnability in Table 1.
Object Model Time Sample
[NPVY24]
$\mathbf{QAC}^{0}$ $n\to m$ Channel,
$a\leq O(\log n)$, Frobenius norm [^1] In this table the Frobenius norm refers to the normalized Frobenius norm.
$2^{n}$ $\mathcal{O}\left(n^{\log^{d}n}\right)$
[wadhwa2024agnostic]
$\mathbf{QAC}^{0}$ $n\to 1$ Channel,
$a\leq\mathcal{O}(\log n)$, Frobenius norm
Agnostic $\mathcal{O}\left(n^{\log^{d}n}\right)$ $\mathcal{O}\left(n^{\log^{d}n}\right)$
[vasconcelos2024learning]
$\mathbf{QAC}^{0}$ Unitary,
$a\leq\mathcal{O}(\log n)$, Frobenius norm
Proper $\mathcal{O}\left(n^{\log^{d}n}\right)$ $\mathcal{O}\left(n^{\log^{d}n}\right)$
[vasconcelos2024learning]
$\mathbf{QAC}^{0}$ Unitary, diamond norm
Improper $\Omega(2^{n})$
[ADOY24]
$\mathbf{QLC}^{0}$ Boolean function
Agnostic $2^{\tilde{\mathcal{O}}\left(n^{1-3^{-d}}\right)}$ $2^{\tilde{\mathcal{O}}\left(n^{1-3^{-d}}\right)}$
This work
$\mathbf{QLC}^{0}$ Boolean function
Agnostic $2^{\tilde{\mathcal{O}}\left(n^{1-2^{-d}}\right)}$ $2^{\tilde{\mathcal{O}}\left(n^{1-2^{-d}}\right)}$
This work
$\mathbf{QLC}^{0}$ $n\to m$ Channel,
Frobenius norm
Agnostic $2^{\tilde{\mathcal{O}}\left(n^{1-2^{-d}}\right)}$ $2^{\tilde{\mathcal{O}}\left(n^{1-2^{-d}}\right)}$
This work
$\mathbf{QLC}^{0}$ $n\to m$ Channel,
spectral norm of Choi repr.
Improper $\Omega(2^{n})$
Table 1: Summary of learning results of shallow circuits with many-qubit gates. Here, $a$ means the number of ancilla. Unless otherwise specified, the column Model refers to the improper standard PAC model. $\mathbf{QLC}^{0}$ means $\mathbf{QAC}^{0}$ with $a=\mathcal{O}(n)$. And we omit dependencies on $\varepsilon,\delta$ and $m$.
Property Testing and Tolerant Testing
Property testing concerns the task of determining whether a given object satisfies a specified property or is far from satisfying it. It is a central area of research in theoretical computer science. Compared with learning, property testing often admits substantially smaller query complexity. In quantum computing, designing efficient testers for various properties has been studied extensively, [PhysRevLett.95.260502, harrow2013testing], including for stabilizer states [10.1145/3717823.3718277], junta states [bao2024learning], quantum junta channels [pmlr-v195-bao23b, chen2024tolerant, bao2025efficient] etc. For an overview, see the survey of Montanaro and de Wolf [montanaro2013survey].
The standard notion of property testing requires an algorithm to distinguish between objects that exactly satisfy the property and those that are far from it. However, this notion lacks robustness to noise. To address this, Parnas, Ron, and Rubinfeld [PARNAS20061012] introduced tolerant property testing, where the goal is to distinguish objects that approximately satisfy the property from those that are far from it. Tolerant testing is particularly significant in the quantum setting, where noise is inherent and unavoidable. Nonetheless, designing efficient tolerant testers is considerably more challenging and, in some cases, provably impossible [10.1145/3618260.3649687].
Tolerant testing is also closely related to agnostic learning. In this work, we propose a quantum tolerant tester for determining whether a quantum channel can be implemented by a $\mathbf{QLC}^{0}$ circuit building on techniques from agnostic learning.
1.1 Our Results
Low-degree Concentration and Hardness
First, similar to [ADOY24], we bound the approximate degree of the output operator of $\mathbf{QLC}^{0}$ channels under low-degree inputs. Specifically, in Section 3, we prove the following theorem:
Theorem 1.1 (informal of Theorem 3.1).
Let $U$ be a depth-$d$ $\mathbf{QLC}^{0}$ circuit working on $n$ inputs qubits. Let $A$ be an operator with degree $\ell$. There exists an operator $M$ with degree $\tilde{\mathcal{O}}\left(n^{1-2^{d}}\ell^{-2^{d}}\right)$ such that
$\left\|UAU^{\dagger}-M\right\|\leq\varepsilon.$
Our proof leverages the unitary dilation technique to ensure that the approximation of large quantum gates is unitary. By embedding the non-unitary, approximated operator into a larger unitary matrix, the technique ensures that the operator can still be managed and analyzed as part of a valid unitary evolution. This allows us to simultaneously apply the low-degree approximation and light-cone argument to a large generalized Toffoli gate. This critical property leads to a tighter bound on the approximate degree than the one established in [ADOY24].
Compared to the $\tilde{\mathcal{O}}\left(n^{1-3^{d}}\ell^{-3^{d}}\right)$ degree upper bound given by [ADOY24], our degree upper bound of $M$ is slightly better. This result directly implies that the Boolean functions computed by $\mathbf{QLC}^{0}$ circuits also have an approximate degree upper bound of $\tilde{\mathcal{O}}\left(n^{1-2^{-d}}\right)$, which matches exactly with the approximate degree upper bound for $\mathbf{LC}^{0}$ circuits, thereby demonstrating the existence of a subexponential time agnostic learning algorithm for $\mathbf{QLC}^{0}$ Boolean functions [doi:10.1137/1.9781611975482.42].
Also, as a direct consequence, we can improve the ancilla lower bound and circuit size lower bound for $\mathbf{QAC}^{0}$ circuits computing the parity function. That is, we have
Corollary 1.2 (informal of Corollary 3.3).
Suppose $U$ is a $\mathbf{QAC}^{0}$ circuit with $n$ input qubits and $a=\tilde{O}\left(n^{1+2^{-d}}\right)$ ancilla qubits. Then $U$ can not approximate the parity function over uniform inputs, with a success probability larger than $1/2+O(d/n)$.
Agnostic Learning
We provide an agnostic learning algorithm for $\mathbf{QLC}^{0}$ channels. The main contribution of our work lies in proposing an algorithmic analysis method based on approximation degree. Thus, in Section 4, we present the following theorem:
Theorem 1.3 (informal of Theorem 4.2).
There is a subexponential time agnostic learning algorithm for $\mathbf{QLC}^{0}$ channels with $n$-qubit input and $\mathrm{polylog}(n)$-qubit outputs with respect to the normalized Frobenius norm.
Our learning algorithm builds upon the framework of Pauli analysis, which involves learning the Pauli coefficients of the Choi state of the quantum channel. Pauli analysis was introduced by Montanaro and Osborne [MO10], and has since become a widely used tool in the design of quantum learning algorithms [doi:10.1137/1.9781611977554.ch43, pmlr-v195-bao23b, NPVY24, arunachalam_et_al:LIPIcs.ICALP.2024.13, Arunachalam2025]. This approach combines with classical shadows [Huang_2020] and Fourier analysis [ODonnell2014], making it particularly effective in scenarios where the Pauli expansion is sparse—such as juntas and low-degree functions.
We further establish hardness for agnostic learning for $\mathbf{QLC}^{0}$ channels under the spectral norm or diamond norm. More specifically, in Section 6, we prove an exponential lower bound on learning $\mathbf{QLC}^{0}$ channels under the spectral norm or diamond norm.
Theorem 1.4 (Hardness of $\mathbf{QAC}^{0}$ channel learning, informal of Theorem 6.1 and Theorem 6.9).
Given an unknown $n$-to-$m$ channel $\Phi\in\mathscr{C}$, learning the Choi representation of $\Phi$ in spectral norm distance requires $\exp\left(\Omega(n)\right)$ queries. Moreover, if $m=\Omega(n)$, then learning $\Phi$ in diamond norm distance requires $\exp\left(\Omega(n)\right)$ queries.
These hardness results also answer a question raised in [wadhwa2024agnostic], demonstrating that agnostic learning of $\mathbf{QLC}^{0}$ channels is impossible under certain stronger norms.
Tolerant Testing
Based on the agnostic learning algorithm for $\mathbf{QLC}^{0}$ channels, we also provide a tolerant testing algorithm for $\mathbf{QLC}^{0}$ channels. Note that a standard learning algorithm is not sufficient to imply this. The tolerant testing problem requires us to determine whether a channel $X$ is sufficiently close to some $\mathbf{QLC}^{0}$ channel or sufficiently far from any $\mathbf{QLC}^{0}$ channel.
Theorem 1.5 (informal of Corollary 5.3).
There exists an $1/\operatorname{poly}(n)$-gap tolerant testing algorithm for $\mathbf{QLC}^{0}$ channels with $n$-qubit input and $\mathrm{polylog}(n)$-qubit outputs. The algorithm has sub-exponential sample and time complexity.
1.2 Proof Overview
Low-degree Concentration
We combine the technique of unitary dilation and operator dilation with the low-degree approximation results from [ADOY24]. The elementary gates in a $\mathbf{QAC}^{0}$ circuit are the multi-qubit CZ-gates, along with all single-qubit unitaries. The single-qubit unitaries are already of degree $1$, so we do not need further action on them. For large CZ-gate acting on $n$ qubits, it can be approximated by an operator with Pauli degree at most $\sqrt{n}$, up to logarithmic factors.
Here comes the dilation technique: The unitary dilation technique allows us to embed an arbitrary operator with bounded spectral norm into a unitary matrix. This is especially useful when handling low-degree approximations of large quantum gates, which lose the property of being a unitary matrix because of the low-degree approximation. We embed such low-degree approximations into larger unitary operators, and use a layer-by-layer argument to prove an approximate-degree upper bound for $\mathbf{QLC}^{0}$ circuits, using similar ideas as in [ADOY24].
The unitary property allows us to have more refined operations when handling a layer of a $\mathbf{QLC}^{0}$ circuit: Specifically, this allows us to combine a light-cone argument with these low-degree operators: For a unitary operator $U_{i}$, we have the identity that $U_{i}^{\dagger}U_{i}=\mathds{1}$. That is, for a layer $U=U_{1}\otimes\cdots\otimes U_{s}$, when we consider the Pauli degree of the operator $U^{\dagger}AU$, with $A$ being a low-degree operator, most of the unitary operators $U_{i}$ will cancel out with themselves. The result is that the Pauli degree will get multiplied by at most a factor of $\max_{i}\deg(U_{i})$. With unitary dilation, we can combine this idea with the low-degree approximation of large CZ-gates, and save a square root factor in this case.
Agnostic Learning for $\mathbf{QLC}^{0}$ circuits in Frobenius norm
Let $\mathscr{C}$ be the set of Choi states of quantum channels implemented by $\mathbf{QLC}^{0}$ circuits. We prove that Algorithm 1 is an agnostic learning algorithm by finding an intermediate class $\mathcal{M}$ such that:
  • Algorithm 1 is an agnostic learning algorithm for the concept class $\mathcal{M}$, and,
  • $\mathscr{C}$ is close to $\mathcal{M}$ in the sense that for any element $C\in\mathscr{C}$, there exists an element $M\in\mathcal{M}$ that is close to $C$.
We employ Pauli analysis (see Section 2.2), and let this intermediate class $\mathcal{M}$ be exactly the operators with a bounded Pauli degree, which we denote as $\mathcal{M}^{\leq d}$ for some $d\in\mathbb{Z}_{\geq 0}$. The advantage of using operators with a low Pauli degree is that it is relatively easy to design agnostic learning algorithms with respect to $M^{\leq d}$: For an arbitrary operator $P$ with Pauli decomposition
$P=\sum_{\sigma\in\left\{0,1,2,3\right\}^{n}}\widehat{P}(\sigma)\mathcal{B}_{\sigma},$
the closest element in $\mathcal{M}^{\leq d}$ in the Frobenius norm is exactly the low-degree part of $P$’s Pauli decomposition
$P^{\leq d}=\sum_{\begin{subarray}{c}\sigma\in\left\{0,1,2,3\right\}^{n}\\ \left|\sigma\right|\leq d\end{subarray}}\widehat{P}(\sigma)\mathcal{B}_{\sigma}.$
So to agnostically learn an arbitrary operator $P$, we only need to learn the low-degree part $P^{\leq d}$, which consists of at most $(3n)^{d}$ Pauli coefficients. This can be achieved efficiently using classical shadow tomography from Huang, Kueng, and Preskill [Huang_2020, Lemma 17].
The remaining task is to prove that any element $C\in\mathscr{C}$ is close to some low-degree operator in $\mathcal{M}=\mathcal{M}^{\leq d}$. This can be achieved by the low-degree concentration results in Theorem 1.1.
Hardness on Learning $\mathbf{QLC}^{0}$ Channels
For $n,a\geq 1$ and a $\mathbf{QAC}^{0}$ circuit $U$, we say $U$ has a clean computation with $a$ ancilla qubits, if there exists a unitary $V$ such that
$U(\left|\varphi\right\rangle\otimes\left|0_{a}\right\rangle)=(V\left|\varphi\right\rangle)\otimes\left|0_{a}\right\rangle.$
In this case we also call $V$ as a $\mathbf{QAC}^{0}$ unitary with $a$ ancilla.
Let $\mathscr{C}$ be the set of Choi representations of $n\to m$ quantum channels implemented by a $\mathbf{QAC}^{0}$ unitary $V$ with $a$ ancilla. We prove that learning the Choi representation $\mathcal{J}(\Phi)\in\mathscr{C}$ under the spectral norm requires an exponential number of queries.
The proof has two steps. In the first step, we prove that learning $n\to 1$ channels requires an exponential number of queries. In the second step, we generalize the result to $n\to m$ channels.
In the first step, we use a reduction from channel hardness to unitary hardness. The main ingredient of the reduction is similar to that in the algorithm [vasconcelos2024learning, Algorithm 1] provided by Vasconcelos and Huang. For a $\mathbf{QAC}^{0}$ unitary $V$ with $a$ ancilla, suppose $V_{i}$ is the circuit related to the $i$-th qubit in the sense that $\mathrm{Tr}_{-i}\>\!\!\left(V\rho V^{\dagger}\right)=\mathrm{Tr}_{-i}\>\!\!\left(V_{i}\rho V_{i}^{\dagger}\right)$. Broadly speaking, the beauty of this algorithm lies in the fact that, using the local information of $V_{i}$, we can sew together the global $V\otimes V^{\dagger}$. Conversely, as a reduction, if learning the global unitary $V\otimes V^{\dagger}$ is difficult, then learning the local $V_{i}$ should also be difficult.
To prove the hardness of learning $\mathbf{QAC}^{0}$ $n\to 1$ channels, we now only need the hardness of learning $V\otimes V^{\dagger}$. This is done with a slightly tailored result from [vasconcelos2024learning, Proposition 7].
In the second step, we embed the $n\to 1$ channels into the $n\to m$ channels. We cannot directly use the partial trace because the spectral norm may increase exponentially under this operation. As an alternative, by imposing constraints on the ancilla size, we accomplish this operation by padding m1m-1 irrelevant qubits.
Acknowledgement
This work was supported by National Natural Science Foundation of China (Grant No. 62332009, and 12347104), Quantum Science and Technology-National Science and Technology Major Project (Grant No. 2021ZD0302901), NSFC/RGC Joint Research Scheme (Grant No. 12461160276), Natural Science Foundation of Jiangsu Province (Grant No. BK20243060) and the New Cornerstone Science Foundation.
1.3 Summary and Future Work
This paper studies agnostic learning for $\mathbf{QAC}^{0}$ circuits and their induced channels. We present a subexponential-time agnostic learning algorithm for linear-size $\mathbf{QAC}^{0}$ $n$-to-$\mathrm{polylog}\left(n\right)$ channels under the normalized Frobenius norm. Furthermore, we establish that learning such channels under the spectral norm or the diamond norm requires exponential queries.
The field of agnostic learning for quantum states and quantum channels is still in its early stage. Here we list several open problems for future research.
  1. 1.
    Is it possible to remove the restriction on the size of the output qubits? Specifically, can we design a subexponential-time agnostic learning algorithm for linear-size $\mathbf{QAC}^{0}$ unitary operators? Although a PAC learning algorithm exists for this setting – with subexponential queries and exponential running time – it remains unclear whether it can be extended to the agnostic case or improved to achieve subexponential runtime.
  2. 2.
    Even subexponential-time algorithms may be too costly for near-term quantum devices. Can we design more efficient agnostic learning algorithms for shallow quantum circuits with geometric structures, such as brickwork or nearest-neighbor architectures?
  3. 3.
    While learning quantum states prepared by shallow circuits has been studied extensively in the PAC setting, little is known about agnostic learning in this context. Recently, the authors in [10.1145/3717823.3718207] proposed an agnostic learning algorithm for product states. A natural question is whether similar techniques can be extended to efficiently learn states generated by $\textsf{state-}\mathbf{QNC}^{0}$ or $\textsf{state-}\mathbf{QLC}^{0}$?
Organization
In Section 2, we will formally introduce agnostic learning and the $\mathbf{QAC}^{0}$ model. In Section 3, we present improved approximation results for $\mathbf{QAC}^{0}$ circuits and their applications. In Section 4, we introduce an agnostic learning algorithm for $\mathbf{QAC}^{0}$ channels with linear ancilla and multiple outputs under the normalized Frobenius norm. In Section 5, we use our agnostic learning algorithm to give a tolerant testing algorithm for $\mathbf{QLC}^{0}$ channels. In Section 6, we propose a reduction for the hardness of quantum channel learning, thereby establishing corresponding hardness results. In Section 1.3, we discuss some future work.
2 Preliminaries
A quantum system $A$ is associated with a finite-dimensional Hilbert space, which we also denote by $A$. The quantum registers in the quantum system $A$ are represented by density operators, which are trace-one positive semi-definite operators, in the Hilbert space $A$. We also use the Dirac notation $\left|\varphi\right\rangle$ to represent a pure state. In this case, we have the convention that $\varphi=\left|\varphi\middle\rangle\middle\langle\varphi\right|$, where here $\varphi$ is a rank-one density operator. For two separate quantum registers $\varphi$ and $\sigma$ from quantum systems $A$ and $B$, The compound register is the Kronecker product $\varphi\otimes\sigma$. A positive operator-valued measure (POVM) is a quantum measurement described by a set of positive semi-definite operators that sum up to the identity. Let $\left\{P_{a}\right\}_{a}$ be a POVM applied on a quantum register $\varphi$, then the probability that the measurement outcome is $a$ is $\mathrm{Tr}\left[P_{a}\varphi\right]$. A quantum channel is a completely positive trace-preserving map. We say quantum channel is an $n\to m$ channel if it takes $n$-qubit as input and outputs $m$ qubits, i.e., $\Phi:\mathcal{M}_{2^{n}}\to\mathcal{M}_{2^{m}}$.
We only consider square matrices in this work. For any integer $n\geq 2$, let $\mathcal{M}_{n}$ be the set of $n\times n$ matrices. The trace of an operator $M\in\mathcal{M}_{n}$ is $\mathrm{Tr}M=\sum_{i=1}^{n}M_{i,i}$ and the normalized trace is $\tau M=\frac{1}{n}\sum_{i=1}^{n}M_{i,i}$. For any matrix $M\in\mathcal{M}_{n}$, we let $\left|M\right|=\sqrt{M^{\dagger}M}$. For any $M,N\in\mathcal{M}_{n}$, the inner product of $M,N$ is $\langle M,N\rangle=\mathrm{Tr}\left[M^{\dagger}N\right]/n$. It is evident that $\left(\mathcal{M}_{n},\langle\cdot,\cdot\rangle\right)$ forms a Hilbert space.
For $p\geq 1$, the normalized Schatten $p$-norm of $M$ is defined to be
$\left\|M\right\|_{p}=\left(\frac{1}{n}\mathrm{Tr}\left[\left|M\right|^{p}\right]\right)^{1/p}.$
For $p=2$, it is not hard to see that $\langle M,M\rangle=\left\|M\right\|_{2}^{2}$. Moreover, $\left\|\cdot\right\|_{p}$ is monotone non-decreasing with respect to $p$ and $\left\|\cdot\right\|_{\infty}=\lim_{p\rightarrow\infty}\left\|\cdot\right\|_{p}$ is the spectral norm. We often use $\left\|A\right\|$ to refer to $\left\|A\right\|_{\infty}$.
Another norm we need to concern is the Frobenius norm
$\left\|M\right\|_{F}=\sqrt{\sum_{i,j}|M_{ij}|^{2}}=\left(\mathrm{Tr}\left[\left|M\right|^{2}\right]\right)^{1/2}.$
which is exactly the Schatten $2$-norm.
The following proposition follows from the sub-additivity of spectral norms.
Proposition 2.1.
For $A=\begin{bmatrix}A_{11}&A_{12}\\ A_{21}&A_{22}\end{bmatrix}$, $\left\|A\right\|\leq\sum_{ij}\left\|A_{ij}\right\|$.
Proposition 2.2 (Hölder inequality).
For operators $A,B$, and $p>1$,
$\left\|AB\right\|_{p}\leq\left\|A\right\|\cdot\left\|B\right\|_{p}$ (1)
Proposition 2.3 ([bhatia2013matrix, Theorem X.1.1]).
For positive semi-definite operators $A,B$ and $r=1/2$,
$\left\|A^{r}-B^{r}\right\|\leq\left\|A-B\right\|^{r}$ (2)
Definition 2.4 (Choi representation and Choi state).
Given a linear map $\Phi:\mathcal{M}_{2^{n}}\to\mathcal{M}_{2^{m}}$, its Choi representation is defined as
$\mathcal{J}(\Phi)=(I\otimes\Phi)(\left|\text{EPR}_{n}\middle\rangle\middle\langle\text{EPR}_{n}\right|)$
where $\left|\text{EPR}_{n}\right\rangle=\sum_{x\in\left\{0,1\right\}^{n}}\left|x\right\rangle\otimes\left|x\right\rangle$. Furthermore, if $\Phi$ is a quantum channel, its Choi state is defined as
$\rho(\Phi)=\frac{1}{2^{n}}\mathcal{J}(\Phi)=\frac{1}{2^{n}}(I\otimes\Phi)(\left|\text{EPR}_{n}\middle\rangle\middle\langle\text{EPR}_{n}\right|).$ (3)
A unitary $U$ induces a channel $\rho\to U\rho U^{\dagger}$, which we refer to as $\Phi_{U}$. We also use $\mathds{1}$ to refer to the identity channel $\Phi_{\mathds{1}}$. For a linear map $\Phi$, its dual map $\Phi^{*}$ satisfies
$\mathrm{Tr}\left[\Phi(X)Y\right]=\mathrm{Tr}\left[X\Phi^{*}(Y)\right].$
Notice that a dual map $\Phi^{*}$ of a quantum channel $\Phi$ is a quantum channel if and only if $\Phi$ is unital, i.e., $\Phi(\mathds{1})=\mathds{1}$. The diamond norm $\left\|\Phi\right\|_{\diamond}$ of an $n$-qubit input map is defined as
$\left\|\Phi\right\|_{\diamond}=\max\left\{\left\|(\Phi\otimes\mathds{1}_{2^{n}})(X)\right\|_{1}:X\in\mathcal{M}_{2^{2n}},\left\|X\right\|_{1}\leq 1\right\}.$ (4)
The following fact holds for unitaries $U$ and $V$,
Proposition 2.5 ([Haah_2023, Proposition 1.6]).
Given unitaries $U$ and $V$, we have
$\left\|\Phi_{U}-\Phi_{V}\right\|_{\diamond}\leq 2\left\|U-V\right\|.$ (5)
For further details regarding distance relations between quantum channels, readers may refer to Yuan and Fung’s work [Yuan_2017].
2.1 Analysis of Boolean Functions
In this subsection, we briefly introduce the theory of analysis of Boolean functions. Readers may refer to O’Donnell’s excellent book [ODonnell2014] for a thorough treatment.
Given a Boolean function $f:\left\{0,1\right\}^{n}\to\mathbb{R}$, its $p$-norm is defined to be $\left\|f\right\|_{p}=\left(\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\mathbf{x}\end{subarray}}\>\!\!\left[\left|f(\mathbf{x})\right|^{p}\right]\right)^{1/p}$ for $p\geq 1$, where $\mathbf{x}$ is a random variable uniformly distributed over $\left\{0,1\right\}^{n}$. Its infinity norm is defined to be $\left\|f\right\|_{\infty}=\lim_{p\to\infty}\left\|f\right\|_{p}=\max_{x}\left|f(x)\right|.$ We will use the notation $\left\|f\right\|=\left\|f\right\|_{\infty}$. Given Boolean functions $f,g:\left\{0,1\right\}^{n}\to\mathbb{R}$, the inner product of $f$ and $g$ is $\langle f,g\rangle=\operatorname*{\mathbb{E}}_{\begin{subarray}{c}\mathbf{x}\end{subarray}}\>\!\!\left[f(\mathbf{x})g(\mathbf{x})\right]$, where $\mathbf{x}$ is uniformly distributed over $\left\{0,1\right\}^{n}$. For any $S\subseteq[n]$, define the Fourier basis $\chi_{S}$ as $\chi_{S}(x)=(-1)^{\sum_{i\in S}x_{i}}$. The set $\left\{\chi_{S}\right\}_{S\subseteq[n]}$ is actually an orthonormal basis for Boolean functions. Any function $f$ admits a Fourier expansion $f=\sum_{S\subseteq[n]}\widehat{f}(S)\chi_{S}$, where $\widehat{f}(S)$ are the Fourier coefficients. Sometimes we consider the case where a Boolean function $f$ has the form $\left\{-1,1\right\}^{n}\to\mathbb{R}$. In this case, $\chi_{S}(x)=\prod_{i\in S}x_{i}$.
The following well-known Parseval’s theorem relates the $2$-norm and Fourier coefficients of a Boolean function.
Theorem 2.6 (Parseval’s theorem).
Let $f:\left\{0,1\right\}^{n}\to\mathbb{R}$ be a Boolean function. Then
$\left\|f\right\|_{2}^{2}=\sum_{S\subseteq[n]}\widehat{f}(S)^{2}.$
For Boolean functions taking values in the set $\left\{-1,1\right\}$, one immediately has $\left\|f\right\|_{2}=1$.
Definition 2.7.
Let $f:\left\{0,1\right\}^{n}\to\mathbb{R}$ be a Boolean function with Fourier expansion
$f=\sum_{S\subseteq[n]}\widehat{f}(S)\chi_{S}.$
Then the degree of $f$ is defined as
$\deg\left(f\right)=\max_{S:\widehat{f}(S)\neq 0}\left|S\right|.$
Definition 2.8 (Approximate Degree).
Let $f:\left\{0,1\right\}^{n}\to\mathbb{R}$ be a Boolean function. For $\varepsilon\in[0,1]$, the approximate degree of $f$ is defined as
$\widetilde{\operatorname{deg}}_{\varepsilon}\left(f\right)=\min_{g:\left\|f-g\right\|\leq\varepsilon}\deg\left(g\right).$
If $\varepsilon$ is not specified, it is $\varepsilon=1/3$.
It is worth noticing that the approximation is with respect to the infinity norm.
2.2 Pauli analysis
Pauli analysis is a generalization of the analysis of Boolean functions to the space of matrices $\mathcal{M}_{2^{n}}$. The Pauli matrices $\mathcal{B}_{0},\dots,\mathcal{B}_{3}$ are
$\mathcal{B}_{0}=\mathds{1}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},\mathcal{B}_{1}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\mathcal{B}_{2}=\begin{bmatrix}0&-i\\ i&0\end{bmatrix},\mathcal{B}_{3}=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$,
which form an orthonormal basis in $\mathcal{M}_{2}$ with respect to the inner product $\langle A,B\rangle=\left(\mathrm{Tr}A^{\dagger}B\right)/2$. For integer $n\geq 1$ and $\sigma\in\left\{0,1,2,3\right\}^{n}$, we define
$\mathcal{B}_{\sigma}=\mathcal{B}_{\sigma_{1}}\otimes\dots\otimes\mathcal{B}_{\sigma_{n}}.$
The set of Pauli matrices $\left\{\mathcal{B}_{\sigma}\right\}_{\sigma\in\left\{0,1,2,3\right\}^{n}}$ forms an orthonormal basis in $\mathcal{M}_{2^{n}}$ with respect to the inner product $\langle A,B\rangle=2^{-n}\mathrm{Tr}A^{\dagger}B$.
For a $2^{n}\times 2^{n}$ matrix $A$, the Pauli expansion of $A$ is
$A=\sum_{\sigma\in\left\{0,1,2,3\right\}^{n}}\widehat{A}(\sigma)\cdot\mathcal{B}_{\sigma},$
where the $\widehat{A}(\sigma)$’s are the Pauli coefficients of $A$. We can then define the degree and the approximate degree of a matrix in a similar manner:
Definition 2.9.
Let $n$ be an integer and $A$ be a $2^{n}\times 2^{n}$ matrix. The degree of $A$ is defined as
$\deg(A)=\max_{\sigma:\widehat{A}(\sigma)\neq 0}\left|\sigma\right|,$
where $\left|\sigma\right|=\left|\left\{i:\sigma_{i}\neq 0\right\}\right|$. For $\varepsilon\in[0,1]$, the approximate degree of $A$ is defined as
$\widetilde{\operatorname{deg}}_{\varepsilon}\left(A\right)=\min_{B:\left\|A-B\right\|\leq\varepsilon}\deg\left(B\right),$
where $\left\|\cdot\right\|$ is the spectral norm.
Let $\mathcal{M}_{n}^{\leq d}$ be the set of $n\times n$ matrices with degree at most $d$. Similar to the classical Parseval’s theorem, one can relate the normalized Schatten $2$-norm with the Pauli coefficients:
Theorem 2.10 (Parseval’s theorem).
Let $A\in\mathcal{M}_{n}$. Then
$\left\|A\right\|_{2}^{2}=\sum_{\sigma\in\left\{0,1,2,3\right\}^{n}}\left|\widehat{A}(\sigma)\right|^{2}.$
We use the notation $A^{\leq k}$ to refer to $\sum_{\sigma:|\operatorname{supp}\left(\sigma\right)|\leq k}\widehat{A}(\sigma)\mathcal{B}_{\sigma}$. The notations $A^{<k},A^{=k},A^{>k},A^{\geq k}$ are similarly defined. From the orthogonality of the Pauli basis and Parseval’s theorem, we know
$\left\|A\right\|_{2}^{2}=\left\|A^{>k}\right\|_{2}^{2}+\left\|A^{\leq k}\right\|_{2}^{2}.$
2.3 Agnostic learning
Kearns, Schapire and Sellie [10.1145/130385.130424] proposed the agnostic PAC learning model, which is a more general learning model, capturing the case where the learning object may not be in the hypothesis class.
Definition 2.11 (Agnostic learning algorithm).
Let $\mathcal{D}$ be a distribution on $\left\{0,1\right\}^{n}\times\left\{0,1\right\}$. For any function $h:\left\{0,1\right\}^{n}\rightarrow\left\{0,1\right\}$, the error of $h$ relative to $\mathcal{D}$ is defined to be $\mathrm{err}_{\mathcal{D}}\left(h\right)=\mathrm{Pr}_{\left(x,y\right)\sim\mathcal{D}}\left[h\left(x\right)\neq y\right]$. Let $\mathscr{C}$ be a concept class, which is a class of functions $c:\left\{0,1\right\}^{n}\to\left\{0,1\right\}$, One defines $\mathrm{opt}(\mathscr{C})=\min_{c\in\mathscr{C}}\mathrm{err}_{\mathcal{D}}\left(c\right)$.
We say that $\mathcal{A}$ is an $(\varepsilon,\delta)$-agnostic learning algorithm for $\mathscr{C}$ if $\mathcal{A}$ has access to the oracle $\mathcal{D}$, and outputs a hypothesis $h$ with probability at least $1-\delta$ that satisfies
$\mathrm{err}_{\mathcal{D}}\left(h\right)\leq\mathrm{opt}(\mathscr{C})+\varepsilon.$
When the hypothesis $h$ we learned satisfies $h\in\mathscr{C}$, we say that the learning algorithm itself is proper.
When the support of the distribution $\mathcal{D}$ satisfies that $\forall(x,y)\in\mathcal{D}$, $y=c(x)$ for some $c\in\mathscr{C}$, it falls back to the standard PAC model. The word “agnostic” comes from the unreliability of the access model.
The agnostic model is believed to be closer to the realistic scenario than the standard PAC model, especially in quantum computing, where noise is unavoidable. However, designing efficient agnostic learning algorithms is generally challenging. Even very few concept classes are known to be agnostically learnable in subexponential time. Bun, Kothari, and Thaler [doi:10.1137/1.9781611975482.42] gave a subexponential time agnostic learning algorithm for the class of functions with approximate degree $n^{c}$ for $c<1$, built on [doi:10.1137/060649057].
Definition 2.12 (Access models).
When querying a quantum channel, we can have different query models. In this work, we consider the following types:
  1. 1.
    Choi State Model. For an $n\to m$ quantum channel $\Phi$, we are given multiple independent copies of the Choi state $\rho(\Phi)$ defined in Definition 2.4, which are quantum states on $n+m$ qubits.
  2. 2.
    Quantum Process Statistical Query (QPSQ) Model [Wadhwa_2025]. For an $n\to m$ quantum channel $\Phi$, in one query we can only obtain classical information about $\Phi$. That is, we can access $\Phi$ by first preparing a quantum state $\varphi$ on $n$ qubits, applying the quantum channel to get the quantum state $\Phi(\varphi)$, and immediately applying a quantum measurement.
In this work, we consider the agnostic learning of quantum channels, which is also referred to as Agnostic Process Tomography introduced by Wadhwa, Lewis, Kashefi and Doosti [wadhwa2024agnostic].
Definition 2.13 (Agnostic Learning of Quantum Channels).
For $\varepsilon,\delta\in(0,1)$, an algorithm $\mathcal{A}$ is an $(\varepsilon,\delta)$-agnostic learner with respect to a set of quantum channels $\mathscr{L}$ and error function $\operatorname{err}$, if given access to an arbitrary channel $\Phi$, the algorithm $\mathcal{A}$ learns a hypothesis $\Phi^{\prime}$ w.p. $1-\delta$ such that
$\operatorname{err}(\Phi,\Phi^{\prime})\leq\min_{\Phi_{C}\in\mathscr{L}}\operatorname{err}(\Phi,\Phi_{C})+\varepsilon.$
When the hypothesis $\Phi^{\prime}$ we learned satisfies $\Phi^{\prime}\in\mathscr{L}$, we say that the learning algorithm is proper. Otherwise, we say that the algorithm is improper.
Remark 2.14.
In the improper learning setting, $\Phi^{\prime}$ is not guaranteed to be a channel.
2.4 Error Functions and Distance
In this work, we assume that the error function $\operatorname{err}$ satisfies the triangle inequality. We mainly focus on the following error distance functions. Given quantum channels $\Phi$ and $\Psi$, we define the following distances.
  • The normalized Frobenius (norm) distance (of Choi representation) is $\left\|\mathcal{J}(\Phi)-\mathcal{J}(\Psi)\right\|_{2}$. This distance can be extended to the improper learning setting. Assuming the learned hypothesis for $\mathcal{J}(\Phi)$ is $M$, the corresponding error distance is $\left\|\mathcal{J}(\Phi)-M\right\|_{2}.$
  • The spectral (norm) distance (of Choi representation) is $\left\|\mathcal{J}(\Phi)-\mathcal{J}(\Psi)\right\|$. This distance can be extended to the improper learning setting. Assuming the learned hypothesis for $\mathcal{J}(\Phi)$ is $M$, the corresponding error distance is $\left\|\mathcal{J}(\Phi)-M\right\|.$
  • The diamond (norm) distance is $\left\|\Phi-\Psi\right\|_{\diamond}$ where the diamond norm is defined in Eq. (4). We do not consider the improper learning setting under this error function.
2.5 Quantum circuit
A quantum circuit involves an input quantum register initialized with some input state $\left|\varphi\right\rangle$, an ancillary quantum register initialized with some fixed state $\left|\psi\right\rangle$, and a series of quantum gates $U_{s},\dots,U_{1}$, where each $U_{i}$ is a unitary operator drawn from a predefined gate set $\mathcal{U}$, and acts on a subset of working qubits. After the computation, the working quantum registers contain the state
$U\left(\left|\varphi\right\rangle\otimes\left|\psi\right\rangle\right)=U_{s}\dots U_{1}\left(\left|\varphi\right\rangle\otimes\left|\psi\right\rangle\right).$
Definition 2.15.
Let $U$ be a unitary implemented by a quantum circuit with $n$ input qubits and $a$ ancilla. Let $\left|\psi\right\rangle$ be a fixed ancilla state on $a$ qubits. We use $\Phi_{k,U,\left|\psi\right\rangle}$ to denote the quantum channel from $n$ qubits to $k$ qubits, implemented by $U$ with ancilla $\left|\psi\right\rangle$, and taking only the first $k$ qubits as output. Formally, for any input state $\varphi$ on $n$ qubits, we have
$\Phi_{k,U,\left|\psi\right\rangle}(\varphi)=\mathrm{Tr}_{\left\{k+1,\dots,n+a\right\}}\>\!\!\left(U\left(\varphi\otimes\left|\psi\middle\rangle\middle\langle\psi\right|\right)U^{\dagger}\right).$
The subscript $k$ is omitted whenever it is clear from the context.
We may get a classical output by applying a computational basis measurement on the first qubit of the output of a quantum circuit. That is, we apply the measurement $\left\{M_{0}=\left|0\middle\rangle\middle\langle 0\right|\otimes\mathds{1},M_{1}=\left|1\middle\rangle\middle\langle 1\right|\otimes\mathds{1}\right\}$. With the input state being $\left|\varphi\right\rangle$ and the ancillae being $\left|\psi\right\rangle$, the probability that we get output $1$ is
$\mathrm{Tr}\left[\left(\left|1\middle\rangle\middle\langle 1\right|\otimes\mathds{1}\right)U\left(\left|\varphi\middle\rangle\middle\langle\varphi\right|\otimes\left|\psi\middle\rangle\middle\langle\psi\right|\right)U^{\dagger}\right].$
We use $C_{U,\left|\psi\right\rangle}$ to denote the above classical output of a quantum channel. When $\left|\psi\right\rangle=\left|0\right\rangle^{a}$ or there is no ancilla, we may simply write $C_{U}$.
In this work, we are concerned with $\mathbf{QAC}^{0}$ circuits, which are polynomial-size constant-depth quantum circuits consisting of single-qubit unitaries and multi-qubit CZ-gates [^2] Some definitions use generalized Toffoli gates. They are equivalent (by the reduction in [10.5555/2011679.2011682]) in our case.. An $n$-qubit CZ-gate is defined as
$\operatorname{CZ}_{n}=\mathds{1}-2\left|1\middle\rangle\middle\langle 1\right|^{\otimes n}.$ (6)
Thus a $\mathbf{QAC}^{0}$ circuit implements a unitary $U=L_{d}M_{d}\dots M_{1}L_{0}$, where $d$ is a constant and each $L_{i}$ is a tensor product of single-qubit unitaries, and each $M_{i}$ is a tensor product of CZ-gates. The depth of this circuit is $d$.
With a slight abuse of notation, we also use $\mathbf{QAC}^{0}$ to represent the family of languages that can be decided by $\mathbf{QAC}^{0}$ quantum circuits. Formally, a language $L$ is in $\mathbf{QAC}^{0}$ if there exists a family of constant-depth and polynomial-size quantum circuits $\left\{C_{n}\right\}_{n\in\mathbb{N}}$ consisting of single-qubit gates and polynomial-size CZ-gates, such that for any $n\in\mathbb{N}$ and $x\in\left\{0,1\right\}^{n}$, if $x\in L$ then $\Pr[C_{n}\left(x\right)=1]\geq 2/3$, and if $x\notin L$, then $\Pr[C_{n}\left(x\right)=0]\geq 2/3$ where $C_{n}(x)$ is the measurement outcome on the output qubits of the circuit $C_{n}$ on input $x$. We say a channel is induced by a $\mathbf{QAC}^{0}$ circuit if the channel can be obtained by implementing a $\mathbf{QAC}^{0}$ circuit and then tracing out part of the qubits.
We also introduce the class of $\mathbf{QLC}^{0}$ circuits, which consists of $\mathbf{QAC}^{0}$ circuits with linear-size ancilla. $\mathbf{QLC}^{0}$ is a quantum counterpart of the classical circuit family $\mathbf{LC}^{0}$, introduced by Chaudhuri and Radhakrishnan [10.1145/237814.237824], which is one of the most interesting subclasses of $\mathbf{AC}^{0}$ and has received significant attention from various perspectives [1663737, 10.1007/978-3-642-11269-0_6, 10.1007/978-3-642-31594-7_65]. A quantum channel induced by a $\mathbf{QLC}^{0}$ circuit is defined analogously.
3 Approximation of $\mathbf{QAC}^{0}$ circuits
In this section, we use unitary dilation combined with the low-degree approximation for CZ-gates, to prove a new low-degree approximation for $\mathbf{QAC}^{0}$ circuits, improving the results of Anshu, Dong, Ou, and Yao [ADOY24].
Theorem 3.1 (Low-degree approximation for $\mathbf{QAC}^{0}$ circuit).
Let $n\geq 1$ be an integer. Let $U$ be an $n$-qubit unitary implemented by a depth-$d$ $\mathbf{QAC}^{0}$ circuit. Then there exist constants $C_{d}$ and $C$ such that the following holds.
For any $2^{n}\times 2^{n}$ operator $A$ with degree at most $\ell$ and $\left\|A\right\|\leq 1$, and any
$r\in\left(2^{9}\log n+C_{d},\left(n/\ell\right)^{3^{-1}\cdot 2^{1-d}}\right)\cup\left(n/\ell,n\right),$ (7)
there exists an operator $M$ such that
$\left\|UAU^{\dagger}-M\right\|\leq dCn\cdot 2^{-2^{-9}r},$ (8)
and
$\deg\left(M\right)\leq\mathcal{O}\left(n^{1-2^{-d}}\cdot\ell^{2^{-d}}\cdot r\right).$
Remark 3.2.
This result is incomparable to [ADOY24]. Although we achieve a better approximation degree upper bound, we have the restriction in Eq. 7 for the error parameter $r$ in this work.
Note that when the degree of $A$ is upper bounded by $\ell=n^{o(1)}$, the above theorem holds for any $r=\operatorname{polylog}(n)$. So we can recover all the results in [ADOY24] with better approximate degree upper bound parameters. For example, Theorem 3.1 implies that any $\mathbf{QAC}^{0}$ circuit that computes $\operatorname{Parity}_{n}$ requires at least $a=\tilde{\Omega}\left(n^{1+2^{-d}}\right)$ ancilla, slightly improving the $\tilde{\Omega}\left(n^{1+3^{-d}}\right)$ lower bound in [ADOY24].
Corollary 3.3.
Suppose $U$ is a $\mathbf{QAC}^{0}$ circuit with depth $d$, and has $n$ input qubits and $a$ ancilla qubits initialized to an arbitrary state. Let $C_{U}:\left\{0,1\right\}^{n}\to\mathbb{R}$ be the function such that $C_{U}(x)$ is the probability that the circuit $U$ output $1$ on input $x$. Then for the parity function defined as
$\operatorname{Parity}_{n}(x)=\bigoplus_{i}x_{i},$
for $a=\tilde{O}\left(n^{1+2^{-d}}\right)$, we have $\Pr\,\>\!\!\left[C_{U}(x)=\operatorname{Parity}_{n}(x)\right]\leq\frac{1}{2}+O(d/n)$, where $x$ is drawn uniformly at random from the set $\left\{0,1\right\}^{n}$.
3.1 Unitary Dilation
We start with the unitary dilation technique used in this work, which is inspired by Montanaro, Shao, and Verdon [montanaro2024lowdegreeapproximationqac0circuits] [^3] This work is withdrawn by the authors due to some flaws in their proof.. The unitary dilation technique allows us to treat non-unitary operators as unitaries by embedding them into larger unitary operators. This is especially useful when applying the polynomial methods [10.1145/28395.28404, tarui1992low] to quantum circuits. In particular, the low-degree approximations of large quantum gates are not necessarily unitary operators. By embedding them into larger unitary operators, we obtain a unitary approximation with a low-degree submatrix.
Definition 3.4 (Unitary Dilation).
Given an operator $A$ with $\left\|A\right\|\leq 1$, its unitary dilation $A^{\uparrow}$ is
$A^{\uparrow}=\begin{bmatrix}A&(\mathds{1}-AA^{\dagger})^{1/2}\\ -(\mathds{1}-A^{\dagger}A)^{1/2}&-A^{\dagger}\end{bmatrix}$. (9)
Specifically, for a unitary operator $V$, we have $V^{\uparrow}=\begin{bmatrix}V&0\\ 0&V^{\dagger}\end{bmatrix}$.
The proof of the following fact is deferred to Appendix A.
Fact 3.5.
For any operator $A$ with $\left\|A\right\|\leq 1$, the operator $A^{\uparrow}$ is a unitary.
Fact 3.6.
For operators $A,B$ with $\left\|A\right\|,\left\|B\right\|\leq 1$ and $\left\|A-B\right\|\leq\varepsilon$, we have
$\left\|A^{\uparrow}-B^{\uparrow}\right\|\leq 5\sqrt{\varepsilon}.$ (10)
Proof.
It suffices for us to bound the spectral norm for each submatrix of $A^{\uparrow}-B^{\uparrow}$ with 2.1.
The top-left and bottom-right parts are simply ABA-B and $(A-B)^{\dagger}$. So they are directly bounded by $\varepsilon$ of the condition.
Since $\left\|A\right\|\leq 1$, we have $(\mathds{1}-AA^{\dagger})^{1/2}\geq 0$. Therefore, we can use 2.3 to deal with the top-right and bottom-left parts:
$\left\|(\mathds{1}-AA^{\dagger})^{1/2}-(\mathds{1}-BB^{\dagger})^{1/2}\right\|\leq\left\|AA^{\dagger}-BB^{\dagger}\right\|^{1/2}.$
Now, by the triangle inequality,
$\left\|AA^{\dagger}-BB^{\dagger}\right\|\leq\left\|A\right\|\cdot\left\|A^{\dagger}-B^{\dagger}\right\|+\left\|B^{\dagger}\right\|\cdot\left\|A-B\right\|\leq 2\varepsilon.$
Finally, with 2.1,
$\left\|A^{\uparrow}-B^{\uparrow}\right\|\leq 2\sqrt{2\varepsilon}+2\varepsilon\leq 5\sqrt{\varepsilon}.$
We also need the notion of operator dilation to incorporate unitary operators obtained from unitary dilation.
Definition 3.7 (Operator Dilation).
Let $n\geq 1$ and $\mathcal{S}=\left\{S_{1},\dots,S_{m}\right\}$ be an ensemble of disjoint subsets of $[n]$. For $\sigma\in\left\{0,1,2,3\right\}^{n}$, the operator dilation for $\mathcal{B}_{\sigma}$ with respect to the ensemble $\mathcal{S}$ is defined as
$\mathcal{B}_{\sigma}^{\Uparrow_{\mathcal{S}}}:=\mathcal{B}_{\sigma}\otimes L(\sigma_{S_{1}})\otimes\cdots\otimes L(\sigma_{S_{m}}),$ (11)
where for any $\tau\in\left\{0,1,2,3\right\}^{*}$,
$L(\tau)=\begin{cases}\left|0\middle\rangle\middle\langle 0\right|&\left|\tau\right|\neq 0,\\ \mathds{1}&\left|\tau\right|=0.\end{cases}$ (12)
For a general operator $A$ with a Pauli expansion $A=\sum_{\sigma}\widehat{A}(\sigma)\mathcal{B}_{\sigma}$, we define
$A^{\Uparrow_{\mathcal{S}}}=\sum_{\sigma}\widehat{A}(\sigma)\mathcal{B}_{\sigma}^{\Uparrow_{\mathcal{S}}}.$ (13)
For simplicity, whenever the ensemble $\mathcal{S}$ is irrelevant or clear from context, $\mathcal{S}$ is omitted, and we simply write $A^{\Uparrow}$.
An interesting observation is that the operator dilation preserves the spectral norm.
Proposition 3.8.
Given a matrix $A\in\mathcal{M}_{2^{n}}$, let $A^{\Uparrow}$ be an operator dilation of $A$. Then,
$\left\|A^{\Uparrow}\right\|=\left\|A\right\|.$ (14)
Proof.
Suppose $A^{\Uparrow}$ is a dilation with respect to the ensemble $\mathcal{S}=\left\{S_{1},\cdots,S_{m}\right\}$. Then $A^{\Uparrow}$ has the following decomposition:
$A^{\Uparrow}$ $=\sum_{T\subseteq[m]}\left|1_{T}0_{T^{c}}\middle\rangle\middle\langle 1_{T}0_{T^{c}}\right|\otimes\sum_{\operatorname{supp}\left(\sigma\right)\subseteq\left(\bigcup_{i\in T}S_{i}\right)^{c}}\widehat{A}(\sigma)B_{\sigma}.$ (15)
Thus,
$\left\|A^{\Uparrow}\right\|=\max_{T}\left\|\sum_{\operatorname{supp}\left(\sigma\right)\subseteq\left(\bigcup_{i\in T}S_{i}\right)^{c}}\widehat{A}(\sigma)B_{\sigma}\right\|=\max_{T}\left\|\tau_{\left[\bigcup_{i\in T}S_{i}\right]}\>\!\!\left(A\right)\right\|.$
Then $\left\|A^{\Uparrow}\right\|\geq\left\|A\right\|$ by setting $T=\emptyset$. Also, by $\left\|\tau_{\left[S\right]}\>\!\!\left(A\right)\right\|\leq\left\|A\right\|$, we conclude that $\left\|A^{\Uparrow}\right\|\leq\left\|A\right\|$. This concludes the proof. ∎
3.2 Proofs of Low-Degree Approximation
In this subsection we prove Theorem 3.1. The main idea we use is that a CZ-gate has a low approximation degree.
Lemma 3.9 (Approximation for CZ-gate, [ADOY24, Corollary 3.3]).
For integer $n\geq 2$ and real number $1<r<n$, there exists an operator $\widetilde{\operatorname{CZ}}_{n}$ such that
$\left\|\operatorname{CZ}_{n}-\widetilde{\operatorname{CZ}}_{n}\right\|\leq 2^{1-2^{-8}r}\log e,\quad\deg\left(\widetilde{\operatorname{CZ}}_{n}\right)\leq\sqrt{nr},\quad\text{and}\quad\left\|\widetilde{\operatorname{CZ}}_{n}\right\|\leq 1.$
With 3.6 and Lemma 3.9, one has the following corollary.
Corollary 3.10 (Dilation of approximation for CZ-gate).
For any $n$-qubit CZ-gate $\operatorname{CZ}_{n}$ and real number $1<r<n$, there exists an operator $\widetilde{\operatorname{CZ}}_{n}$ such that
$\left\|\operatorname{CZ}_{n}^{\uparrow}-\widetilde{\operatorname{CZ}}_{n}^{\uparrow}\right\|\leq\sqrt{10}\cdot 2^{-2^{-9}r}\log e$ (16)
and
$\deg\left(\widetilde{\operatorname{CZ}}_{n}\right)\leq\sqrt{nr}.$
We begin by proving low-degree approximations for a single layer of a $\mathbf{QAC}^{0}$ circuit.
Lemma 3.11.
There exists an absolute constant $C>0$ such that the following holds:
Let $n\geq 1$ be an integer and $U=\bigotimes_{i}\operatorname{CZ}_{S_{i}}$ be a layer of CZ-gates, where $S_{1},\dots,S_{m}$ are disjoint subsets of $[n]$. For each $i$, $\operatorname{CZ}_{S_{i}}$ is a $\operatorname{CZ}_{\left|S_{i}\right|}$ gate acting on qubits in $S_{i}$. For any integer $\ell$ and $r\in(1,\sqrt{n/\ell})\cup(n/\ell,n)$ and $2^{n}\times 2^{n}$ operator $A$ with degree at most $\ell$, there exists an operator $M$ such that
$\left\|UAU^{\dagger}-M\right\|\leq Cn\cdot 2^{-2^{-9}r}\cdot\left\|A\right\|$ (17)
and
$\deg\left(M\right)\leq 4n^{1/2}\ell^{1/2}r^{1/2}.$ (18)
The key idea of the proof is as follows. Let $t$ be a parameter to be specified later. Divide the CZ-gates into three parts:
  • For gates of size $\leq t$, we use a lightcone argument.
  • For gates of size between $t$ and $t^{2}$, we perform a low-degree approximation and use a lightcone argument. To incorporate the lightcone argument with low-degree approximations, we need to use the unitary dilation and operator dilation techniques explained earlier.
  • For gates of size $t^{2}$, we perform a low-degree approximation and use a direct degree argument.
The complete proof is deferred to Appendix A.
We note that, in contrast to [ADOY24], our approximation requires a range constraint on the parameter $r$. Fortunately, if we require an inverse-polynomial error, we only need $r=O(\log n)$ and this remains within the applicable scope of the lemma.
Single-qubit unitaries do not change the degree of an operator. So, after applying Lemma 3.11 for $d$ times, we obtain Theorem 3.1, which we rephrase below for the reader’s convenience.
See 3.1
Proof.
Let $p=\tilde{C}n\cdot 2^{-2^{-9}r}$ where $\tilde{C}$ is the constant in Lemma 3.11. Choose $C=2\tilde{C}$ and $C_{d}$ large enough such that $\tilde{C}2^{-2^{-9}C_{d}}\leq 2^{1/d}-1$. The constant selection ensures that $(1+p)^{d}\leq 2$.
Let $U$ be $U=V_{d}L_{d}V_{d-1}L_{d-1}\cdots V_{1}L_{1}$ where $V$ stands for multi-qubit unitary layer and $L$ stands for single-qubit unitary layer. Write $U_{\leq 0}=U_{0}=\mathds{1},U_{i}=V_{i}L_{i}$ and $U_{\leq i}=U_{i}U_{\leq i-1}$.
We define operators $\left\{M_{i}\right\}$ inductively for $i=0,\dots,d$. Set $M_{0}=A$.
We will prove by induction that for any $i=0,\dots,d-1$, we have
$\left\|U_{\leq i}AU_{\leq i}^{\dagger}-M_{i}\right\|\leq iCn\cdot 2^{-2^{-9}r}$ (19)
and
$\left\|M_{i}\right\|\leq(1+p)^{i},\quad\deg\left(M_{i}\right)\leq\mathcal{O}\left(n^{1-2^{-i}}\ell^{2^{-i}}r\right).$ (20)
The base case is $i=0$. Indeed, for Eq. 19 it is easy to check that
$\left\|U_{\leq 0}AU_{\leq 0}^{\dagger}-M_{0}\right\|=\left\|A-A\right\|=0.$
and Eq. 20 holds for $i=0$. Now suppose Eq. 19 and Eq. 20 hold for some $i$. We carefully choose $r$ such that Lemma 3.11 remains applicable to the parameters $A\leftarrow M_{i}$ and $U\leftarrow U_{i}$, and one can find $M_{i+1}$ such that
$\left\|U_{i}M_{i}U_{i}^{\dagger}-M_{i+1}\right\|\leq\tilde{C}n2^{-2^{-9}r}\cdot\left\|M_{i}\right\|$ (21)
and
$\deg\left(M_{i+1}\right)\leq\mathcal{O}\left(\left(n^{1-2^{-{i}}}\ell^{2^{-{i}}}r\right)^{1/2}r^{1/2}\right)=\mathcal{O}\left(n^{1-2^{-{i+1}}}\ell^{2^{-{i+1}}}r\right).$ (22)
Recall that $\left\|M_{i}\right\|=(1+p)^{i}\leq(1+p)^{d}\leq 2$. Using the triangle inequality, we now bound the spectral norm:
$\left\|U_{\leq i+1}AU_{\leq i+1}^{\dagger}-M_{i+1}\right\|$ $=\left\|U_{\leq i}AU_{\leq i}^{\dagger}-U_{i+1}^{\dagger}M_{i+1}U_{i+1}\right\|$
$\leq\left\|U_{\leq i}AU_{\leq i}^{\dagger}-M_{i}\right\|+\left\|M_{i}-U_{i+1}^{\dagger}M_{i+1}U_{i+1}\right\|$
$\leq iCn2^{-2^{-9}r}+2\tilde{C}n2^{-2^{-9}r}=(i+1)Cn2^{-2^{-9}r}.$
We finish our induction process by estimating the norm of $M_{i+1}$.
$\;\;\left\|U_{i}M_{i}U_{i}^{\dagger}-M_{i+1}\right\|\leq\tilde{C}n2^{-2^{-9}r}\cdot\left\|M_{i}\right\|$
$\Rightarrow\left\|M_{i+1}\right\|\leq(1+p)\cdot\left\|M_{i}\right\|\leq(1+p)^{i+1}.$
Taking $M$ as $M_{d}$ completes the proof. ∎
4 Agnostic Learning for $\mathbf{QLC}^{0}$ channels
In this section, we present an agnostic learning algorithm for $\mathbf{QLC}^{0}$ channels with respect to the Frobenius norm as in Algorithm 1. This algorithm employs the classical shadow tomography introduced by Huang, Kueng and Preskill [Huang_2020].
Lemma 4.1 (Classical Shadows for Low-degree Pauli Operators, [Huang_2020, Lemma 17]).
For any $\varepsilon,\delta\in(0,1)$, there exists an algorithm such that:
Given a quantum state $\rho$ and a list of observables $\left\{O_{1},\cdots,O_{M}\right\}$, satisfying $\left\|O_{i}\right\|\leq 1$ for all $i\in[M]$, the algorithm outputs estimates $\left\{\hat{o}_{i}\right\}$ satisfying
$\left|\hat{o}_{i}-\mathrm{Tr}\left(O_{i}\rho\right)\right|\leq\varepsilon,\forall i\in[M]$ (23)
with probability at least $1-\delta$.
When the $O_{i}$ are degree-$d$ Pauli operators, one needs
$N=\mathcal{O}\left(\frac{3^{d}\log(M/\delta)}{\varepsilon^{2}}\right)$
copies of $\rho$, and the (classical) computation time is $\mathcal{O}(dNM).$
Parameters : $d\in\mathcal{Z}_{\geq 0}$, and $\varepsilon,\delta\in(0,1)$.
Input : Access to the Choi state $\rho(\Phi)$ of an $n\to m$ quantum channel $\Phi$, where the Choi state of a quantum channel is defined in Definition 2.4
Output : An approximation $L$ such that $\left\|L-\left(\mathcal{J}(\Phi)\right)^{\leq d}\right\|_{2}\leq\varepsilon$ with probability $1-\delta$
1:Perform the classical shadow tomography algorithm in Lemma 4.1 with parameters $\varepsilon\leftarrow\frac{\varepsilon^{2}}{d(n+m)^{d}2^{m}}$ and $\delta\leftarrow\delta$, for the quantum state $\rho(\Phi)$ with Pauli observables $\left\{\mathcal{B}_{\sigma}:\left|\sigma\right|\leq d\right\}$, to get a list of estimates $\left\{\hat{o}_{\sigma}:\left|\sigma\right|\leq d\right\}$.
2:For each $\sigma$ with $\left|\sigma\right|\leq d$, set $\hat{\alpha}_{\sigma}=\hat{o}_{\sigma}\cdot 2^{-m}$.
3:Output $L=\sum_{\sigma:\left|\sigma\right|\leq d}\hat{\alpha}_{\sigma}\mathcal{B}_{\sigma}$.
Algorithm 1 Channel-Learning($\Phi$, $d$, $\delta$, $\varepsilon$)
Our main result shows that for any $\varepsilon\geq 1/\operatorname{poly}(n)$ and $\delta\in(0,1)$, Algorithm 1 is an $(\varepsilon,\delta)$-agnostic learner with respect to the class of quantum channels implemented by $\mathbf{QLC}^{0}$ circuits, using only sub-exponential queries. Formally,
Theorem 4.2.
Given $0<\varepsilon,\delta<1$ and $n,m,d,a\geq 1$, and assume $\mathscr{C}$ is a set of Choi representations of $n\to m$ quantum channels implemented by depth-$d$ $\mathbf{QAC}^{0}$ circuits with $a=\mathrm{poly}(n)$ ancilla. For any $\varepsilon$ satisfying
$\varepsilon\geq-\exp\left(n^{2^{-d-2}}/m\right),$
let
$D=\tilde{\mathcal{O}}\left((n+a)^{1-2^{-d}}\cdot m^{2^{-d}+1}\cdot\log^{2}(1/\varepsilon)\right),$
where $\tilde{\mathcal{O}}$ hides log factors. Algorithm 1 with parameters $(d,\varepsilon,\delta)\leftarrow(D,\varepsilon,\delta)$ is an $(\varepsilon,\delta)$-agnostic learner with respect to $\mathscr{C}$ and error function $\operatorname{err}(X,Y)=\left\|X-Y\right\|_{2}$. The number of queries to $\Phi$ and the running time of the algorithm are both
$\mathcal{O}\left(\operatorname{poly}(n,m)\cdot 4^{m}\cdot(n+m)^{3D}\cdot\frac{\log(1/\delta)}{\varepsilon^{4}}\right).$
In particular, when $m=\mathrm{polylog}(n),\varepsilon=1/\mathrm{poly}(n)$, and $a=\mathcal{O}(n)$, the number of queries to $\Phi$ and the running time are both $2^{\tilde{O}\left(n^{1-2^{-d}}\right)}$.
We can also get a variant of Theorem 4.2 where we don’t have the restriction on $\varepsilon\geq 1/\operatorname{poly}(n)$. For technical reasons, however, we need to slightly increase the number of queries.
Theorem 4.3.
Given $0<\varepsilon,\delta<1$ and $n,m,d,a\geq 1$, and assume $\mathscr{C}$ is a set of Choi representations of $n\to m$ quantum channels implemented by depth-$d$ $\mathbf{QAC}^{0}$ circuits with $a=\mathrm{poly}(n)$ ancilla. For
$D=\tilde{\mathcal{O}}\left((n+a)^{1-3^{-d}}\cdot m^{3^{-d}+1/2}\cdot\log(1/\varepsilon)\right),$
Algorithm 1 with parameters $(d,\varepsilon,\delta)\leftarrow(D,\varepsilon,\delta)$, is an $(\varepsilon,\delta)$-agnostic learner with respect to $\mathscr{C}$ and error function $\operatorname{err}(X,Y)=\left\|X-Y\right\|_{2}$. The number of queries to $\Phi$ and the running time are both
$\mathcal{O}\left(\operatorname{poly}(n,m)\cdot 4^{m}\cdot(n+m)^{3D}\cdot\frac{\log(1/\delta)}{\varepsilon^{5}}\right).$
In particular, when $m=\mathrm{polylog}(n),\varepsilon=1/\mathrm{poly}(n)$, and $a=\mathcal{O}(n)$, the number of queries to $\Phi$ and the running time are both $2^{\tilde{O}\left(n^{1-3^{-d}}\right)}.$
Note that the above theorems assume the Choi state query model. While the Choi-state access model is analytically convenient, it isn’t easy to implement in real physical environments, where a more natural model is the Quantum Process Statistical Query (QPSQ) model. [^4] See Definition 2.12 for the definitions of these query models. Luckily, in this model we can use [NPVY24, Lemma 40] to achieve the same effect of the shadow tomography algorithm in Lemma 4.1, with a $(n+m)^{D}$ number of queries overhead. Thus, in the QPSQ model, we also have an agnostic learning algorithm, but at the cost of increasing the number of samples by a factor of $(n+m)^{D}$.
Corollary 4.4.
Given $0<\varepsilon,\delta<1$, $n,m,d\geq 1$ and $a\geq 0$, and assume $\mathscr{C}$ is a set of Choi representations of $n\to m$ quantum channels implemented by depth-$d$ $\mathbf{QAC}^{0}$ circuits with $a$ ancilla. For
$D=\tilde{\mathcal{O}}\left((n+a)^{1-3^{-d}}\cdot m^{3^{-d}+1/2}\cdot\log(1/\varepsilon)\right),$
Under the QPSQ access model, there is an $(\varepsilon,\delta)$-agnostic learner with respect to $\mathscr{C}$ and error function $\operatorname{err}(X,Y)=\left\|X-Y\right\|_{2}$, The number of queries to $\Phi$ and the computation time are both
$\mathcal{O}\left(\operatorname{poly}(n,m)\cdot 4^{m}\cdot(n+m)^{4D}\cdot\frac{\log(1/\delta)}{\varepsilon^{5}}\right).$
4.1 Analysis of Agnostic Learning Algorithm
We start by analyzing the algorithm in Theorem 4.2. The proof of Theorem 4.3 is similar, where we will simply replace 4.7 with 4.11.
We prove that Algorithm 1 is an agnostic learning algorithm by finding an intermediate set of objects $\mathcal{M}$ that is close to the concept class $\mathscr{C}$, and at the same time, Algorithm 1 is also an agnostic learning algorithm with respect to $\mathcal{M}$. In our work, this intermediate set $\mathcal{M}$ is the low-degree operators.
Lemma 4.5 (Skeleton of Agnostic Learning).
Given $0<\varepsilon_{1},\varepsilon_{2},\delta<1$ and $n,m\geq 1$, and assume $\mathscr{C}$ is a set of Choi representations of $n\to m$ channels, and $\mathcal{A}$ is a learning algorithm. Recall $\mathcal{M}_{2^{n+m}}^{\leq d}$ is the set of matrices with Pauli degree at most $d$. For simplicity, denote $\mathcal{M}^{\leq d}=\mathcal{M}_{2^{n+m}}^{\leq d}$. Suppose the following two conditions hold for some error function $\operatorname{err}$ on matrices satisfying the triangle inequality.
  1. (1)
    Algorithm $\mathcal{A}$ is an agnostic learning algorithm with respect to the concept class $\mathcal{M}^{\leq d}$: There exists $\varepsilon_{1}>0$ such that, given access to any channel $\Phi$, the algorithm $\mathcal{A}$ outputs an $L\in\mathcal{M}_{2^{n+m}}$ in time $T(n,\delta,\varepsilon_{1})$ such that with probability at least $1-\delta$,
    $\operatorname{err}(\mathcal{J}(\Phi),L)\leq\min_{M\in\mathcal{M}^{\leq d}}\operatorname{err}(\mathcal{J}(\Phi),M)+\varepsilon_{1}.$ (24)
  2. (2)
    The class $\mathscr{C}$ is close to $\mathcal{M}^{\leq d}$: Any Choi representation in $\mathscr{C}$ can be approximated by some matrix in $\mathcal{M}^{\leq d}$. That is, there exists $\varepsilon_{2}>0$ such that for any $\mathcal{J}(\Phi)\in\mathscr{C}$,
    $\min_{M\in\mathcal{M}^{\leq d}}\operatorname{err}(\mathcal{J}(\Phi),M)\leq\varepsilon_{2}.$ (25)
Then $\mathcal{A}$ is an $(\varepsilon_{1}+\varepsilon_{2},\delta)$-agnostic learner for $\mathscr{C}$ in time $T(n,\delta,\varepsilon_{1})$ with respect to error function $\operatorname{err}$.
Proof.
Fix any $n\to m$ channel $\Phi$ to be learned. Let $\mathcal{J}(\Psi^{*})\in\mathscr{C}$ be the minimizer of
$\min_{\mathcal{J}(\Psi)\in\mathscr{C}}\operatorname{err}(\mathcal{J}(\Phi),\mathcal{J}(\Psi)).$
Let $\text{opt}=\operatorname{err}(\mathcal{J}(\Phi),\mathcal{J}(\Psi^{*}))$. By the triangle inequality and then Eq. 25,
$\min_{M\in\mathcal{M}^{\leq d}}\operatorname{err}(\mathcal{J}(\Phi),M)$ $\leq\operatorname{err}(\mathcal{J}(\Phi),\mathcal{J}(\Psi^{*}))+\min_{M\in\mathcal{M}^{\leq d}}\operatorname{err}(\mathcal{J}(\Psi^{*}),M)$
$\leq\text{opt}+\varepsilon_{2}.$
Suppose algorithm $\mathcal{A}$ outputs $\tilde{M}$ as an approximation of $\mathcal{J}(\Phi)$. By Eq. 24,
[…]
$\leq\text{opt}+\varepsilon_{1}+\varepsilon_{2}.$
From now on, choose $\operatorname{err}(X,Y)=\left\|X-Y\right\|_{2}$. For Algorithm 1, we verify the two conditions in Lemma 4.5, separately by Lemma 4.6 and 4.7.
We first check the condition (1) that Algorithm 1 is an agnostic learning algorithm with respect to the class $\mathcal{M}^{\leq d}$.
Lemma 4.6.
For any parameters $1\leq d\leq n+m$, and $0<\varepsilon,\delta<1$, Algorithm 1 satisfies condition (1) in Lemma 4.5 with $\varepsilon_{1}\leftarrow\varepsilon$. That is, given access to an $n\to m$ channel $\Phi$, Algorithm 1 outputs an $L\in\mathcal{M}_{2^{n+m}}$ such that with probability $1-\delta$,
$\left\|L-\mathcal{J}(\Phi)\right\|_{2}\leq\min_{M\in\mathcal{M}_{2^{n+m}}^{\leq d}}\left\|M-\mathcal{J}(\Phi)\right\|+\varepsilon.$ (26)
Moreover, Algorithm 1 uses
$\mathcal{O}\left(\frac{4^{m}(n+m)^{2d}\log\left(\frac{(n+m)^{d}}{\delta}\right)}{\varepsilon^{4}}\right)$
queries to $\rho(\Phi)$ and the computational time is
$\mathcal{O}\left(\frac{4^{m}d^{3}(n+m)^{3d}\log\left(\frac{(n+m)^{d}}{\delta}\right)}{\varepsilon^{4}}\right).$
Proof of Lemma 4.6.
Estimating Pauli coefficients of the Choi representation can be finished by applying classical shadow to the Choi state. This fact also indicates that in Algorithm 1, classical shadow tomography is essentially estimating the Pauli coefficients.
$\mathrm{Tr}\left[\rho(\Phi)\mathcal{B}_{\sigma}\right]=\frac{1}{2^{n}}\cdot 2^{n+m}\cdot\widehat{\mathcal{J}(\Phi)}(\sigma)=2^{m}\widehat{\mathcal{J}(\Phi)}(\sigma).$ (27)
For simplicity, define $\alpha_{\sigma}=\widehat{\mathcal{J}(\Phi)}(\sigma)$. Now, using Theorem 4.1 with argument $\tilde{\varepsilon}=\frac{\varepsilon^{2}}{d(n+m)^{d}2^{m}}$, we know that the approximation results $\left\{\hat{\alpha}_{\sigma}\right\}$ satisfy
$\left|\hat{\alpha}_{\sigma}-\alpha_{\sigma}\right|\leq\tilde{\varepsilon}$.
Due to Theorem 2.10,
$\left\|L-\mathcal{J}(\Phi)^{\leq d}\right\|_{2}\leq\sqrt{d(n+m)^{d}2^{m}\tilde{\varepsilon}}=\varepsilon.$ (28)
Let $M^{*}\in\mathcal{M}_{2^{n+m}}^{\leq d}$ be the matrix minimizing $\left\|\mathcal{J}(\Phi)-M^{*}\right\|_{2}$, in fact, we have
$M^{*}=\sum_{\sigma:|\operatorname{supp}\left(\sigma\right)|\leq d}\alpha_{\sigma}\mathcal{B}_{\sigma}=\mathcal{J}(\Phi)^{\leq d}.$
Thus,
$\left\|\mathcal{J}(\Phi)-L\right\|_{2}$ $\leq\left\|\mathcal{J}(\Phi)-M^{*}\right\|_{2}+\left\|L-M^{*}\right\|_{2}$
$\leq\left\|\mathcal{J}(\Phi)-M^{*}\right\|_{2}+\varepsilon$
$=\min_{M\in\mathcal{M}_{2^{n+m}}^{\leq d}}\left\|\mathcal{J}(\Phi)-M\right\|_{2}+\varepsilon.$
Now, let us turn to the task of verifying condition (2) in Lemma 4.5. We use the following proposition, which is based on the low-degree approximation results of $\mathbf{QLC}^{0}$ circuits.
Proposition 4.7.
Given $0<\varepsilon<1$, $n,m,d\geq 1$ satisfying $m\cdot\log(1/\varepsilon)<n^{2^{-d-2}}$, $a\geq 0$, and assume $\mathscr{C}$ is a set of Choi representations of $n\to m$ quantum channels implemented by depth-$d$ $\mathbf{QAC}^{0}$ circuits with $a=\mathrm{poly}(n)$ ancilla initialized in a pure state $\psi$. For any $D$ satisfying
$D\leq\tilde{\mathcal{O}}\left((n+a)^{1-2^{-d}}m^{1+2^{-(d-1)}}\log^{2}(1/\varepsilon)\right),$
the class $\mathscr{C}$ satisfies condition (2) in Lemma 4.5 with $d\leftarrow D$ and $\varepsilon_{2}\leftarrow\varepsilon$. That is, for any Choi representation $\mathcal{J}(\Phi)\in\mathscr{C}$, we have
$\min_{M\in\mathcal{M}^{\leq D}}\left\|M-\mathcal{J}(\Phi)\right\|_{2}\leq\varepsilon.$ (29)
In fact, we have the stronger conclusion that Eq. 29 holds within the spectral distance. That is, in the approximate degree parlance, for any $\mathcal{J}(\Phi)\in\mathscr{C}$, we have
$\widetilde{\operatorname{deg}}_{\varepsilon}\left(\mathcal{J}(\Phi)\right)\leq\tilde{\mathcal{O}}\left((n+a)^{1-2^{-d}}m^{1+2^{-(d-1)}}\log^{2}(1/\varepsilon)\right).$
To prove 4.7, this work requires some results in [ADOY24]. We therefore reformulate the results accordingly.
Lemma 4.8 (Approximate degree for state, [Anshu2023concentrationbounds], [ADOY24, Corollary 3.2]).
Let $\left|\psi\right\rangle$ be an $\ell$-qubit pure state. Then for any $r\in(\sqrt{n},n)$ and $\varepsilon=2^{-\frac{r^{2}}{2^{8}n}}$. It holds that
$\widetilde{\operatorname{deg}}_{\varepsilon}\left(\left|\psi\middle\rangle\middle\langle\psi\right|^{\otimes n}\right)\leq lr$
We only care about the EPR state with the following corollary.
Corollary 4.9.
For any integer $n>0$ and $0<\varepsilon<2^{-1/2^{8}}$,
$\widetilde{\operatorname{deg}}_{\varepsilon}\left(2^{-n}\left|\text{EPR}_{n}\middle\rangle\middle\langle\text{EPR}_{n}\right|\right)\leq\mathcal{O}\left(\sqrt{n\cdot\log(1/\varepsilon)}\right)$
Note that when $\varepsilon<2^{-n}$, the inequality holds trivially since $\deg\left(2^{-n}\left|\text{EPR}_{n}\middle\rangle\middle\langle\text{EPR}_{n}\right|\right)\leq 2n$.
Lemma 4.10 ([ADOY24, Lemma 2.12]).
Given integers $m\leq n$ and $0<\varepsilon<1$, it holds that for $M\in\mathcal{M}_{2^{n}}$ and any $2^{m}\times 2^{m}$ density operator $\varphi$,
$\widetilde{\operatorname{deg}}_{\varepsilon}\left(\mathrm{Tr}_{n-m+1,\ldots,n}\>\!\!\left((\mathds{1}\otimes\varphi)M\right)\right)\leq\widetilde{\operatorname{deg}}_{\varepsilon}\left(M\right)$
Proof of 4.7.
With [ADOY24, Fact 7.1],
$2^{-m}\mathcal{J}(\Phi)=\left\langle\psi\right|\left(\mathds{1}\otimes U^{T}\right)\left(2^{-m}\left|\text{EPR}_{m}\middle\rangle\middle\langle\text{EPR}_{m}\right|\otimes\mathds{1}_{n+a-m}\right)\left(\mathds{1}\otimes\bar{U}\right)\left|\psi\right\rangle.$
Let $\varepsilon^{\prime}=\varepsilon/3$. Apply Corollary 4.9 with parameter $\varepsilon\leftarrow\varepsilon^{\prime}$, we get that there is an operator $M$ with degree $\ell=\mathcal{O}\left(\sqrt{m\log(1/\varepsilon)}\right)$ such that
$\left\|M-\left(2^{-m}\left|\text{EPR}_{m}\middle\rangle\middle\langle\text{EPR}_{m}\right|\otimes\mathds{1}\right)\right\|\leq\varepsilon^{\prime}.$
Then, apply Theorem 3.1 with parameters $r\leftarrow\tilde{\mathcal{O}}(\log(1/\varepsilon^{\prime}))$ and $A\leftarrow M$, where the condition $m\cdot\log(1/\varepsilon)\leq n^{C_{d}}$ ensures that the condition of the Theorem 3.1 holds. There exists an operator $\tilde{M}$ with degree
$\mathcal{O}\left((n+a)^{1-2^{-d}}m^{2^{-d}}\log^{1+2^{-d}}(1/\varepsilon)\right)$
such that
$\left\|\tilde{M}-\left(\mathds{1}\otimes U^{T}\right)M\left(\mathds{1}\otimes\bar{U}\right)\right\|\leq\varepsilon^{\prime}.$
Now,
$\;\;\left\|\tilde{M}-\left(\mathds{1}\otimes U^{T}\right)\left(2^{-m}\left|\text{EPR}_{m}\middle\rangle\middle\langle\text{EPR}_{m}\right|\otimes\mathds{1}\right)\left(\mathds{1}\otimes\bar{U}\right)\right\|$
$\leq\left\|\tilde{M}-\left(\mathds{1}\otimes U^{T}\right)M\left(\mathds{1}\otimes\bar{U}\right)\right\|+\left\|\left(\mathds{1}\otimes U^{T}\right)\left(M-2^{-m}\left|\text{EPR}_{m}\middle\rangle\middle\langle\text{EPR}_{m}\right|\right)\left(\mathds{1}\otimes\bar{U}\right)\right\|$
$\leq\varepsilon^{\prime}+2\varepsilon^{\prime}=\varepsilon.$
As a conclusion,
$\;\;\widetilde{\operatorname{deg}}_{\varepsilon}\left(\left(\mathds{1}\otimes U^{T}\right)\left(2^{-m}\left|\text{EPR}_{m}\middle\rangle\middle\langle\text{EPR}_{m}\right|\otimes\mathds{1}_{n+a-m}\right)\left(\mathds{1}\otimes\bar{U}\right)\right)$
$=\tilde{\mathcal{O}}\left((n+a)^{1-2^{-d}}m^{2^{-d}}\log^{1+2^{-d}}(1/\varepsilon)\right).$
With Lemma 4.10,
$\widetilde{\operatorname{deg}}_{\varepsilon}\left(2^{-m}\mathcal{J}(\mathcal{E}_{m,U,\psi})\right)\leq\widetilde{\operatorname{deg}}_{\varepsilon}\left(\left(\mathds{1}\otimes U^{T}\right)\left(2^{-m}\left|\text{EPR}_{m}\middle\rangle\middle\langle\text{EPR}_{m}\right|\otimes\mathds{1}_{n+a-m}\right)\left(\mathds{1}\otimes\bar{U}\right)\right).$
Reorganizing it as
$\widetilde{\operatorname{deg}}_{\varepsilon}\left(\mathcal{J}(\mathcal{E}_{m,U,\psi})\right)\leq\tilde{\mathcal{O}}\left((n+a)^{1-2^{-d}}m^{1+2^{-(d-1)}}\log^{2}(1/\varepsilon)\right).$
The proof completes here. ∎
We are now ready to prove Theorem 4.2.
Proof of Theorem 4.2.
Plugging 4.7 and Lemma 4.6 into Lemma 4.5, we conclude that Algorithm 1 is indeed an agnostic learning algorithm for $\mathbf{QLC}^{0}$ channels, thus proving Theorem 4.2. ∎
To prove Theorem 4.3, we use the following proposition in place of 4.7.
Proposition 4.11.
Given $0<\varepsilon<1$, $n,m,d\geq 1$ and $a\geq 0$, and assume $\mathscr{C}$ is a set of Choi representations of $n\to m$ quantum channels implemented by depth-$d$ $\mathbf{QAC}^{0}$ circuits with $a=\mathrm{poly}(n)$ ancilla initialized in a pure state $\psi$. For any $D$ satisfying
$D\leq\tilde{\mathcal{O}}\left((n+a)^{1-3^{-d}}m^{3^{-d}+1/2}\log(1/\varepsilon)\right),$
the class $\mathscr{C}$ satisfies condition (2) in Lemma 4.5 with $d\leftarrow D$ and $\varepsilon_{2}\leftarrow\varepsilon$. I.e., for any $\mathcal{J}(\Phi)\in\mathscr{C}$, we have
$\min_{M\in\mathcal{M}^{\leq D}}\left\|M-\mathcal{J}(\Phi)\right\|_{2}\leq\varepsilon.$ (30)
The proof of this proposition is highly similar to that of 4.7, and the argument is completed by simply replacing Theorem 3.1 with [ADOY24, Lemma 3.5].
5 Tolerant Testing for $\mathbf{QLC}^{0}$ channels
Here, we exhibit an application of our agnostic learning algorithm in tolerant testing.
Definition 5.1 ($\varepsilon$-gap Tolerant testing for channels).
Let $0<\varepsilon_{1}<\varepsilon_{2}<1$ such that $\varepsilon_{2}-\varepsilon_{1}\geq\varepsilon$. Let $\Phi$ be an unknown channel, and $\mathscr{C}$ a set of Choi representations of quantum channels. Assuming that one of the following two cases is true, the task is to determine which one.
  1. 1.
    $\min_{C\in\mathscr{C}}\left\|C-\mathcal{J}(\Phi)\right\|_{2}\leq\varepsilon_{1}$.
  2. 2.
    $\min_{C\in\mathscr{C}}\left\|C-\mathcal{J}(\Phi)\right\|_{2}\geq\varepsilon_{2}$.
By augmenting the agnostic learning algorithm with an estimator for the distance between the target channel, we can directly obtain a tolerant testing algorithm.
Proposition 5.2.
Suppose $\mathcal{A}$ is an $(\varepsilon,\delta)$-agnostic learning algorithm for $\mathscr{C}$ under normalized Frobenius norm, then there is a $2\varepsilon$-gap tolerant testing algorithm with the same sample and computational complexity.
Proof.
Suppose $\mathcal{A}$ outputs $L$, then one just checks whether
$\left\|L-\mathcal{J}(\Phi)\right\|_{2}\geq\frac{\varepsilon_{1}+\varepsilon_{2}}{2}$ (31)
or not. For the yes case, determine that it falls under the case $\geq\varepsilon_{2}$. Conversely, determine that it falls under the case $\leq\varepsilon_{1}$. The property of agnostic learning
$\left\|L-\mathcal{J}(\Phi)\right\|_{2}\leq\min_{C\in\mathscr{C}}\left\|\mathcal{J}(C)-\mathcal{J}(\Phi)\right\|_{2}+\varepsilon$ (32)
ensures correctness. ∎
By slightly decreasing the value of $\varepsilon$, the algorithm can further accommodate errors in distance computation. In particular, by using low-degree operators as a bridge, we can efficiently compute the distance, leading to the following corollary:
Corollary 5.3.
Assume $\mathscr{C}$ is the Choi representation of a family of $\mathbf{QLC}^{0}$ $n\to\operatorname{polylog}(n)$ channels. There exists an $1/\operatorname{poly}(n)$-gap tolerant testing algorithm with sub-exponential sample and time complexity.
Proof.
By applying the agnostic learning algorithm in Algorithm 1, we obtain an $L$. Recall that $\left\|L-\mathcal{J}(\Phi)^{\leq d}\right\|_{2}\leq\varepsilon$, after estimating $\left\|\mathcal{J}(\Phi)\right\|_{2}$ and obtaining a value $v$ such that $\left|v-\left\|\mathcal{J}(\Phi)\right\|_{2}\right|\leq\varepsilon$, we have
$\left\|L-\mathcal{J}(\Phi)\right\|_{2}$ $\leq\left\|L-\mathcal{J}(\Phi)^{\leq d}\right\|_{2}+\left\|\mathcal{J}(\Phi)-\mathcal{J}(\Phi)^{\leq d}\right\|_{2}$
$=\varepsilon+\sqrt{\left\|\mathcal{J}(\Phi)\right\|_{2}^{2}-\left\|\mathcal{J}(\Phi)^{\leq d}\right\|_{2}^{2}}$
$\leq\varepsilon+\sqrt{(v+\varepsilon)^{2}-\left(\left\|L\right\|_{2}-\varepsilon\right)^{2}}$
$\leq\sqrt{v^{2}-\left\|L\right\|_{2}^{2}}+2\sqrt{\varepsilon}+\varepsilon$
$\leq\sqrt{v^{2}-\left\|L\right\|_{2}^{2}}+3\sqrt{\varepsilon}.$
And conversely, since $L$ is low-degree,
$\left\|L-\mathcal{J}(\Phi)\right\|_{2}$ $\geq\left\|\mathcal{J}(\Phi)-\mathcal{J}(\Phi)^{\leq d}\right\|_{2}$
$=\sqrt{\left\|\mathcal{J}(\Phi)\right\|_{2}^{2}-\left\|\mathcal{J}(\Phi)^{\leq d}\right\|_{2}^{2}}$
$\geq\sqrt{(v-\varepsilon)^{2}-\left(\left\|L\right\|_{2}+\varepsilon\right)^{2}}$
$\geq\sqrt{v^{2}-\left\|L\right\|_{2}^{2}}-2\sqrt{\varepsilon}.$
Still, since $L$ is low-degree, $\left\|L\right\|_{2}$ is computable. Thus, if we apply agnostic learning algorithm for $\varepsilon^{\prime}=o\left((\varepsilon_{2}-\varepsilon_{1})^{2}\right)$, we have accomplished the estimation of the distance with error $o\left(\varepsilon_{2}-\varepsilon_{1}\right)$, thereby completing the tolerant testing . ∎
Note that if we can approximate the value of $\left\|\mathcal{J}(\Phi)^{\leq d}\right\|_{2}$ more efficiently (without requiring knowledge of its approximate form), we could derive a more efficient testing algorithm.
6 Hardness on Learning $\mathbf{QAC}^{0}$ Channels
In this section, we prove our hardness results of learning quantum channels in terms of the spectral norm.
Theorem 6.1 (Hardness of $\mathbf{QAC}^{0}$ channel learning within spectral norm).
Given $n,m,a,d\geq 1$. Assume $\mathscr{C}$ is the set of all Choi representations of $n\to m$ channels implemented by a depth-$d$ $\mathbf{QAC}^{0}$ unitaries $U$ with $a$ ancilla, where $a\geq m$. Then, given an unknown channel $\mathcal{J}(\Phi)\in\mathscr{C}$, learning $\mathcal{J}(\Phi)$ up to a spectral distance $\frac{1}{n}$ with probability at least $1-\frac{1}{n^{2}}$ requires $\exp\left(\Omega(n)\right)$ queries.
This hardness is proved via a reduction to unitary hardness. In Section 6.1, we present the reduction, and the proof is given in Section 6.2.
6.1 Hardness Reduction
We now introduce the key lemma to be used in the reduction, which enables separate processing of each output qubit and ultimately synthesizes the target unitary matrix.
Definition 6.2 (Local inversion).
$V_{i}$ is a local inversion of $U$ for the $i$-th qubit if and only if
$UV_{i}=U_{-i}\otimes\mathds{1}_{i}$
for some unitary $U_{-i}$ acting only on $\left\{i\right\}^{c}$.
Lemma 6.3 ([HLB+24, Eq. (5)]).
Given a unitary $U\in\mathcal{M}_{n}$, for each $i\in[n]$, suppose $V_{i}$ is a local inversion of $U$ for the $i$-th qubit, then
$U\otimes U^{\dagger}=S\cdot\prod_{i=1}^{n}\left(V_{i}\cdot S_{i}\cdot V_{i}^{\dagger}\right)$ (33)
where $S_{i}$ is the SWAP operator for $i$-th and $(i+n)$-th qubit, $S=\prod_{i=1}^{n}S_{i}$ and $V_{i}\cdot S_{i}\cdot V_{i}^{\dagger}$ are the abbreviations of $(V_{i}\otimes\mathds{1})\cdot S_{i}\cdot(V_{i}^{\dagger}\otimes\mathds{1})$.
The following lemma allows us to focus solely on estimating Heisenberg-evolved Pauli observables.
Lemma 6.4.
For $x\in\left\{0,1,2,3\right\}$ and $i\in[n]$, let $\mathcal{B}_{x}^{(i)}$ be an $n$-qubit operator that applies the Pauli operator $\mathcal{B}_{x}$ on $i$-th qubit and $\mathds{1}$ on other qubits. That is
$\mathcal{B}_{x}^{(i)}=\mathds{1}^{\otimes i-1}\otimes\mathcal{B}_{x}\otimes\mathds{1}^{\otimes n-i-1}.$
Then,
$V_{i}S_{i}V_{i}^{\dagger}=\frac{1}{2}\sum_{x\in\left\{0,1,2,3\right\}}V_{i}\mathcal{B}_{x}^{(i)}V_{i}^{\dagger}\otimes\mathcal{B}_{x}^{(i+n)}$ (34)
where $S_{i}$ is the SWAP operator for $i$-th and $(i+n)$-th qubits.
Remark 6.5.
$V_{i}\mathcal{B}_{x}^{(i)}V_{i}^{\dagger}$ are called the Heisenberg-evolved Pauli observables.
Proof.
With the fact $S_{i}=\frac{1}{2}\sum_{x\in\left\{0,1,2,3\right\}}\left(\mathcal{B}_{x}^{(i)}\otimes\mathcal{B}_{x}^{(i+n)}\right)$,
$V_{i}S_{i}V_{i}^{\dagger}=\frac{1}{2}\sum_{x\in\left\{0,1,2,3\right\}}(V_{i}\otimes\mathds{1})\left(\mathcal{B}_{x}^{(i)}\otimes\mathcal{B}_{x}^{(i+n)}\right)(V_{i}^{\dagger}\otimes\mathds{1})=\frac{1}{2}\sum_{x\in\left\{0,1,2,3\right\}}V_{i}\mathcal{B}_{x}^{(i)}V_{i}^{\dagger}\otimes\mathcal{B}_{x}^{(i+n)}.$
We then present the reduction directly in Reduction 2. The key intuition is that to learn Heisenberg-evolved Pauli observables, we can attempt to learn the channel: $\rho\to V_{i}\left(\rho\otimes\mathds{1}\right)V_{i}^{\dagger}$. Notably, the dual of this channel precisely corresponds to the sub-channel that restricts the circuit output to the $i$-th qubit.
Input : Given the Choi state of an $n$-input $\mathbf{QAC}^{0}$ quantum channel $\Phi_{U}$, and access to an algorithm $\mathcal{A}$ which outputs an approximation of a Choi representation when given access to the Choi state, and a real number $0<\delta<1$.
Output : An approximation $Q\approx U\otimes U^{\dagger}$ w.p. $1-\delta$.
1:Repeat the following for $i\in\left\{1,2,\cdots,n\right\}$;
  1. A
    Let $V_{i}$ be a local inversion of $V$ corresponding to the $i$-th qubit and $\Phi_{V_{i}}(\rho)=\mathrm{Tr}_{-i}\>\!\!\left(V_{i}\rho V_{i}^{\dagger}\right)$ be the $n\to 1$ channel with only the $i$-th qubit reserved. Given Choi state $\rho(\Phi_{U})$, trace out all but the $i$-th qubit to get $\rho(\Phi_{V_{i}})$.
  2. B
    Learn an approximation $M_{i}$ for $\mathcal{J}(\Phi_{V_{i}})$ using algorithm $\mathcal{A}$ with access to $\rho(\Phi_{V_{i}})$ and failure probability $\delta/n$.
  3. C
    Learn $Q_{i,x}$ as an approximation of Heisenberg-evolved Pauli observables $V_{i}\mathcal{B}_{x}^{(i)}V_{i}^{\dagger}$ (defined in Remark 6.5): set $Q_{i,0}=\mathds{1}$ and $Q_{i,x}=\mathrm{Tr}_{1}\>\!\!\left(M_{i}^{T}(\mathcal{B}_{x}\otimes\mathds{1})\right)$ for $x=1,2,3$.
  4. D
    Sewing the local inversion $V_{i}S_{i}V_{i}^{\dagger}$ from Heisenberg-evolved Pauli observables: set $Q_{i}=\sum_{x}Q_{i,x}\otimes\mathcal{B}_{x}^{(i+n)}$.
2:Return $Q=S\cdot\prod_{i=1}^{n}Q_{i}$.
Reduction 2 Unitary-Reduction($\Phi_{U},\mathcal{A}$, $\delta$)
Unlike [vasconcelos2024learning] and [HLB+24], we perform no operations during step B of our protocol.
In other words, this reduction transfers unitary hardness to channel hardness.
Theorem 6.6.
Given an integer $n>0$ and real numbers $0<\varepsilon,\delta<1$, let $\mathcal{U}$ be a set of $n$-qubit unitaries. Suppose given access to any channel $\Phi$ of the form $\mathrm{Tr}_{[n-1]}\>\!\!\left(\Phi_{U}\right)$ with single qubit output where $U\in\mathcal{U}$, algorithm $\mathcal{A}$ outputs a hypothesis $H$ w.p. at least $1-\delta$ such that
$\left\|H-\mathcal{J}(\Phi)\right\|\leq\varepsilon.$ (35)
Then, there exists an algorithm $\mathcal{A^{\prime}}$, given access to a channel $\Phi_{U}$ with $U\in\mathcal{U}$, algorithm $\mathcal{A^{\prime}}$ performs Reduction 2 using $\mathcal{A}$ as an oracle in step B, then outputs a hypothesis $H^{\prime}$ w.p. at least $1-n\delta$ such that
$\left\|H^{\prime}-U\otimes U^{\dagger}\right\|\leq 9\cdot n\varepsilon.$ (36)
Moreover, if $\mathcal{A}$ has sample complexity $N(\varepsilon,\delta,n)$, then $\mathcal{A}^{\prime}$ has sample complexity
$N^{\prime}(\varepsilon,\delta,n)=n\cdot N(\varepsilon/9n,\delta/n,n).$
The above theorem essentially demonstrates that if the unitaries of a circuit class are hard to learn, then their corresponding channels should also exhibit learning hardness.
Proof of Theorem 6.6.
Let $V$ denote the implicitly declared unitary matrix, $V_{i}$ be the local inversion (Definition 6.2) of $V$ corresponding to the $i$-th qubit and $\Phi_{V_{i}}$ be the $n\to 1$ channel with only the $i$-th qubit reserved.
To begin with the analysis of the reduction, we assume that in step B, one can learn $Q_{i,x}$ such that
$\left\|M_{i}-J(\Phi_{V_{i}})\right\|\leq\varepsilon.$
And starting from now on, we fix $\varepsilon$ and define $\varepsilon_{V}$ which will subsequently be derived from $\varepsilon$.
Now, assuming that step 1 of the reduction has been successfully completed, by a hybrid argument, with a straightforward calculation we can show that under the condition where step 1 provides a valid estimation, $Q$ is indeed a valid approximation of $U\otimes U^{\dagger}$. Formally, suppose $\left\|Q_{i}-V_{i}S_{i}V_{i}^{\dagger}\right\|\leq\varepsilon_{V}<1/n$, then
$\left\|Q-U\otimes U^{\dagger}\right\|\leq 3n\varepsilon_{V}.$ (37)
We defer the proof to the Appendix A.
The remaining task is to justify the correctness of step 1, which is also the most intriguing part.
We first prove that in step A, performing the partial trace on the Choi representation $\mathcal{J}(\Phi_{U})$ indeed yields the Choi representation $\mathcal{J}(\Phi_{V_{i}})$ corresponding to the sub-channel acting only on the $i$-th qubit. This is divided into two steps: the first step verifies the relationship between $\Phi_{U}$ and $\Phi_{V}$.
$\mathcal{J}(\Phi_{U})$ $=(\mathds{1}\otimes\left\langle 0_{a}\right|)\left((\mathds{1}\otimes U)\left|\text{EPR}_{n}\middle\rangle\middle\langle\text{EPR}_{n}\right|(\mathds{1}\otimes U^{\dagger})\right)(\mathds{1}\otimes\left|0_{a}\right\rangle)$
$=(\mathds{1}\otimes V)\left|\text{EPR}_{n}\middle\rangle\middle\langle\text{EPR}_{n}\right|(\mathds{1}\otimes V^{\dagger})$
$=\mathcal{J}(\Phi_{V})$
The second equality follows from the clean computation assumption.
The second step verifies that the partial trace of the Choi state indeed yields the Choi state corresponding to the sub-channel, since
$\mathrm{Tr}_{-i}\>\!\!\left(\mathcal{J}(\Phi_{V})\right)$ $=\mathrm{Tr}_{-i}\>\!\!\left((\mathds{1}\otimes U^{\prime}\otimes\mathds{1}_{i})(\mathds{1}\otimes V_{i})\left|\text{EPR}_{n}\middle\rangle\middle\langle\text{EPR}_{n}\right|(\mathds{1}\otimes V_{i})^{\dagger}(\mathds{1}\otimes U^{\prime}\otimes\mathds{1}_{i})^{\dagger}\right)$
$=\mathrm{Tr}_{-i}\>\!\!\left((\mathds{1}\otimes V_{i})\left|\text{EPR}_{n}\middle\rangle\middle\langle\text{EPR}_{n}\right|(\mathds{1}\otimes V_{i})^{\dagger}\right)$
$=\mathcal{J}(\Phi_{V_{i}}).$
From the hypothesis, in step B, we have
$\left\|M_{i}-\mathcal{J}(\Phi_{V_{i}})\right\|\leq\varepsilon.$
Let us focus on the step C. We review some facts originating from quantum information theory (see the book of Watrous [watrous2018theory] for more information): if an $n\to 1$ channel $\Phi$ has the form $\rho\to\mathrm{Tr}_{-i}\>\!\!\left(V\rho V^{\dagger}\right)$, then its dual channel $\Phi^{*}$ will take the form $\rho_{i}\to V^{\dagger}(\rho_{i}\otimes\mathds{1}_{-i})V$. Moreover, the Choi representation of the dual channel and that of the original channel satisfy the following relation:
$\mathcal{J}(\Phi)^{T}=\mathcal{J}(\Phi^{*})$
Thus,
$\left\|M_{i}^{T}-\mathcal{J}(\Phi_{V_{i}}^{*})\right\|=\left\|M_{i}-\mathcal{J}(\Phi_{V_{i}})\right\|=\left\|M_{i}^{T}-\mathcal{J}(\Phi_{V_{i}})^{T}\right\|\leq\varepsilon.$
That is, $M_{i}^{T}$ is actually the approximation to the Choi representation of $\rho_{i}\to V^{\dagger}(\rho_{i}\otimes\mathds{1}_{-i})V$. This channel is of importance, since when inputting $\mathcal{B}_{x}$, the output corresponds precisely to the desired Heisenberg-evolved Pauli observables. By the fact
$\Phi(X)=\mathrm{Tr}_{\text{input}}\>\!\!\left(\mathcal{J}(\Phi)(X^{T}\otimes\mathds{1})\right),$ (38)
we have:
$\left\|Q_{i,x}-V_{i}\mathcal{B}_{x}^{(i)}V_{i}^{\dagger}\right\|$ $=\left\|\mathrm{Tr}_{1}\>\!\!\left(M_{i}^{T}(\mathcal{B}_{x}\otimes\mathds{1})\right)-\mathrm{Tr}_{1}\>\!\!\left(\mathcal{J}(\Phi_{i}^{*})(\mathcal{B}_{x}\otimes\mathds{1})\right)\right\|$
$\leq 2\left\|M_{i}^{T}(\mathcal{B}_{x}\otimes\mathds{1})-\mathcal{J}(\Phi_{i}^{*})(\mathcal{B}_{x}\otimes\mathds{1})\right\|$
$\leq 2\left\|M_{i}^{T}-\mathcal{J}(\Phi_{i}^{*})\right\|\cdot\left\|\mathcal{B}_{x}\otimes\mathds{1}\right\|$
$\leq 2\varepsilon.$
Finally, for step D, with Lemma 6.4,
$\left\|Q_{i}-V_{i}S_{i}V_{i}^{\dagger}\right\|$ $\leq\frac{1}{2}\sum\left\|Q_{i,x}\otimes\mathcal{B}_{x}^{(i+n)}-V_{i}\left(\mathcal{B}_{x}^{(i)}\otimes\mathcal{B}_{x}^{(i+n)}\right)V_{i}^{\dagger}\right\|$
$\leq\frac{1}{2}\sum\left\|Q_{i,x}-V_{i}\mathcal{B}_{x}^{(i)}V_{i}^{\dagger}\right\|\cdot\left\|\mathcal{B}_{x}^{(i+n)}\right\|$
$\leq\frac{1}{2}\cdot 3\cdot 2\varepsilon=3\varepsilon$
The final constant is 3, since we can always output perfect replicas on the $\mathds{1}$ case. Ultimately, we obtain $\varepsilon_{V}=3\varepsilon$ and the total error distance $9\cdot n\varepsilon$.
This completes the proof of Theorem 6.6. ∎
6.2 Hardness Results
Let us briefly recall the background of the reduction, where we take the same setting as in Vasconcelos and Huang’s work [vasconcelos2024learning]. Suppose $U$ is a unitary generated by a circuit with $(n+a)$ qubits. We only consider circuits where the ancilla qubits are initialized to the $\left|0_{a}\right\rangle$ state and the computation is clean. Since the computation is clean, $U$ implicitly characterizes a unitary matrix $V$ acting on $n$ input qubits.
$U\left(\left|\varphi\right\rangle\otimes\left|0^{a}\right\rangle\right)=\left(V\left|\varphi\right\rangle\right)\otimes\left|0^{a}\right\rangle.$ (39)
While it may appear that learning a unitary with ancilla is equivalent to the no ancilla case, it is crucial to note that introducing ancilla changes the properties of the unitary (like degree). We may use $\mathbf{QAC}^{0}$ unitary $U$ with $a$ ancilla to refer to the implicit unitary $V$.
Now we are ready to prove Theorem 6.1. Let us first examine the learning hardness of $\mathbf{QAC}^{0}$ unitary. We use a result of Vasconcelos and Huang [vasconcelos2024learning], tailored to our needs. For completeness, we provide a proof.
Lemma 6.7 (Tailored from [vasconcelos2024learning, Propsition 7]).
Consider an unknown $n$-qubit unitary $U$ generated by a depth-$d$ $\mathbf{QAC}^{0}$ circuit with $a$ ancilla. Then, distinguishing whether $U\otimes U^{\dagger}$ equals the $\mathds{1}$ or is $\frac{1}{3}$-far from $\mathds{1}$ in diamond distance with probability $2/3$ requires $\exp\left(\Omega(n)\right)$ queries.
Proof.
For $x,y\in\left\{0,1\right\}^{n}$, let $U_{x}$ be the unitary,
$U_{x}\left|y\right\rangle=(-1)^{\delta_{xy}+1}\left|y\right\rangle.$ (40)
which can be constructed as
$U_{x}=\mathcal{B}_{\bar{x}}\operatorname{CZ}_{n}\mathcal{B}_{\bar{x}}.$
$U_{x}$ is in the depth-$d$ $\mathbf{QAC}^{0}$ circuit with ancilla even for $a=0$ and $d=1$. However, distinguishing $\mathds{1}$ from one of $U_{x}$ has a Grover Search lower bound from Bennett, Bernstein, Brassard and Vazirani’s work [doi:10.1137/S0097539796300933] with queries at least $\exp\left(\Omega(n)\right)$. This also means distinguishing $\mathds{1}$ from one of $U_{x}\otimes U_{x}^{\dagger}$ has an exponential bound. ∎
Combining this hardness assumption for $\mathbf{QAC}^{0}$ unitary with Theorem 6.6, we can establish hardness results even for channels with minimal input requirements.
Lemma 6.8 (Hardness of $\mathbf{QAC}^{0}$ $n\to 1$ channel learning with spectral distance).
Given integers $a\geq 0$ and $n,d\geq 1$. Assuming $\mathscr{C}$ is the set of Choi representations of $n\to 1$ channels induced by depth-$d$ $\mathbf{QAC}^{0}$ unitaries $U$ with $a$ ancilla. Then, given an unknown Choi $\mathcal{J}(\Phi)\in\mathscr{C}$, learning $\mathcal{J}(\Phi)$ up to a spectral distance $\frac{1}{n}$ with probability at least $1-\frac{1}{n^{2}}$ requires $\exp\left(\Omega(n)\right)$ queries.
Proof.
One minor technical detail to note is that after performing the reduction in Theorem 6.6, we obtain a hypothesis $H$ such that
$\left\|H-U\otimes U^{\dagger}\right\|\leq 1/12.$
To relate this result to the diamond norm, we locally compute the nearest unitary matrix $\tilde{U}$ of $H$ in the spectral norm (though this may require exponential computation time, it does not affect sample complexity). At this point, we have
$\left\|U\otimes U^{\dagger}-\tilde{U}\right\|\leq\left\|U\otimes U^{\dagger}-H\right\|+\left\|H-\tilde{U}\right\|\leq 1/6.$
Since the small spectral norm distance between unitary matrices leads to a small diamond norm distance (2.5), this completes the proof with
$\left\|\Phi_{U\otimes U^{\dagger}}-\Phi_{\tilde{U}}\right\|_{\diamond}\leq 2\left\|U\otimes U^{\dagger}-\tilde{U}\right\|\leq 1/3.$
We now begin the proof of Theorem 6.1.
Proof of Theorem 6.1.
Suppose $\mathcal{S}$ is the set of channels induced by depth-$1$ $\mathbf{QAC}^{0}$ circuits with $n$ inputs, sharing no ancilla and taking the first qubit as output. From Lemma 6.8, we know that learning $\mathcal{S}$ requires exponential samples.
Fix a channel $\Phi\in\mathcal{S}$, let $\tilde{\Phi}(\rho)=\Phi(\rho)\otimes\left|0_{m-1}\middle\rangle\middle\langle 0_{m-1}\right|$. Since $a\geq m$, $\tilde{\Phi}$ can be implemented by a depth-$d$ $\mathbf{QAC}^{0}$ circuit with $n$ inputs and $a$ ancilla.
Since
$\mathcal{J}(\tilde{\Phi})=\mathcal{J}(\Phi)\otimes\left|0_{m-1}\middle\rangle\middle\langle 0_{m-1}\right|,$
we now have
[…]
Thus, if we are able to have $\left\|M-\mathcal{J}(\tilde{\Phi})\right\|\leq 1/n$, by taking $\tilde{M}=(\mathds{1}\otimes\left\langle 0_{m-1}\right|)M(\mathds{1}\otimes\left|0_{m-1}\right\rangle)$, we have $\left\|\tilde{M}-\mathcal{J}(\Phi)\right\|\leq 1/n$. Learning $\mathcal{S}$ needs exponential queries, therefore, learning $\mathbf{QAC}^{0}$ channels given in the condition also needs exponential queries.
This completes the proof of Theorem 6.1. ∎
The spectral norm of the Choi representations is almost $2^{m}$. However, our hardness result only demonstrates that approximation is hard within a distance of $1/n$. When the number of outputs is sufficiently large (like $\sqrt{n}$), compared to $2^{m}$, the value $1/n$ is too small to be an appropriate error. This inspires us to propose the other hardness result in the diamond norm, which is powerful with a large number of output qubits.
Theorem 6.9 (Hardness of $\mathbf{QAC}^{0}$ channel learning within diamond norm distance).
Given integers $n,m\geq 1$ and $a,d\geq 0$. Assuming $\mathscr{L}$ is the set of $n\to m$ channels induced by depth-$d$ $\mathbf{QAC}^{0}$ circuits with $a$ ancilla. Then, given an unknown channel $\Phi\in\mathscr{L}$, learning $\Phi$ up to diamond norm distance $1/3$ with probability $2/3$ requires $\exp\left(\Omega(m)\right)$ queries.
Proof.
Notice that the lower bound in Lemma 6.7 relies on a class of hard unitary $\left\{U_{n}\right\}$ that can be implemented by $\mathbf{QAC}^{0}$ circuits. We can directly embed a size $k$ unitary $U_{k}$ into a channel with $k$ outputs, thereby deriving corresponding lower bounds for $\mathbf{QAC}^{0}$ channels within diamond norm distance. ∎
6.3 Hardness of Finding the Nearest Low-Degree Operator
In this subsection, we present the hardness results further reduced from the aforementioned hardness of $\mathbf{QAC}^{0}$ channel learning.
We recall the following proposition: $\mathbf{QLC}^{0}$ circuits all admit low-degree approximations. Consequently, for a given quantum channel, finding its closest low-degree operator remains computationally hard.
Theorem 6.10 (Hardness of finding the nearest low-degree operator).
Assume $M\in\mathcal{M}_{2^{n}}$. For degree $D=\Omega\left(\sqrt{n}\cdot\operatorname{poly}(\log n)\right)$, finding
$\operatorname{argmin}_{X\in\mathcal{M}_{2^{n}}^{\leq D}}\left\|X-M\right\|$ (41)
up to a spectral distance $\frac{1}{n}$ with probability at least $1-\frac{1}{n^{2}}$ requires $\exp\left(\Omega(n)\right)$ queries.
Proof.
We focus on the case where the $\mathbf{QAC}^{0}$ channel $\Phi_{U}$ has at most depth $1$, has zero ancilla and only outputs the first qubit. With Proposition 4.7,
$\widetilde{\operatorname{deg}}_{1/n}\left(\mathcal{J}(\Phi_{U})\right)=\sqrt{n}\cdot\textrm{poly}(\log n),$
i.e.,
$\exists X\in\mathcal{M}_{2^{n}}^{\leq D},\left\|X-\mathcal{J}(\Phi_{U})\right\|\leq\frac{1}{n}.$
Assuming the hypothesis of this theorem holds, the closest low-degree operator to $\mathcal{J}(\Phi_{U})$ would yield a $1/n$-approximation to $\mathcal{J}(\Phi_{U})$, thereby contradicting Lemma 6.8. ∎
Appendix A Proofs
Proof of 3.5.
Let $D_{A}=\sqrt{I-A^{\dagger}A}$ and $U=A^{\uparrow}$, then
$U=\begin{bmatrix}A&D_{A^{\dagger}}\\ D_{A}&-A^{\dagger}\end{bmatrix}$.
With a direct calculation,
$U^{\dagger}U=\begin{bmatrix}A^{\dagger}A+D_{A}D_{A}&A^{\dagger}D_{A^{\dagger}}-D_{A}A^{\dagger}\\ D_{A^{\dagger}}A-AD_{A}&D_{A^{\dagger}}D_{A^{\dagger}}+AA^{\dagger}\end{bmatrix}=\begin{bmatrix}\mathds{1}&A^{\dagger}D_{A^{\dagger}}-D_{A}A^{\dagger}\\ D_{A^{\dagger}}A-AD_{A}&\mathds{1}\end{bmatrix}$.
This shows that we only need to prove $A^{\dagger}D_{A^{\dagger}}-D_{A}A^{\dagger}=D_{A^{\dagger}}A-AD_{A}=0$. By symmetry, it suffices to prove $D_{A^{\dagger}}A-AD_{A}=0$. Define $B=D_{A}^{2}$ and $C=D_{A^{\dagger}}^{2}$. Now, we use an inductive argument to prove that for all $n\geq 1$, we have
$AB^{n}=C^{n}A.$
For the base case where $n=1$, it is to check that $AB=CA=A-AA^{\dagger}A$. Now for any $n\geq 2$, suppose $AB^{n-1}=C^{n-1}A$ and $AB=CA$. Then,
$AB^{n}=(AB^{n-1})B=(C^{n-1}A)B=C^{n-1}(AB)=C^{n-1}CA=C^{n}A.$
Thus, by a limit argument, for all elementary function $f:\mathbb{R}\to\mathbb{R}$, we have
$Af(B)=f(C)A.$
In particular, for $f=\sqrt{x}$, we conclude that $D_{A^{\dagger}}A-AD_{A}=0$.
Proof of Lemma 3.11.
The proof is similar to the one in [ADOY24].
When $r>n/\ell$, we have $n^{1/2}\ell^{1/2}r^{1/2}\geq n$. So choosing $M=UAU^{\dagger}$ directly completes the proof. In the remaining proof, we can assume $r\leq\sqrt{n/\ell}$.
Recall that $\operatorname{CZ}_{S_{i}}$ acts on $n_{i}=\left|S_{i}\right|$ qubits on the set $S_{i}\in[n]$. We divide the CZ-gates into three parts with parameter $t=\sqrt{n/\ell}>r$. Let $T_{0}=\left\{i:n_{i}\leq t\right\},T_{1}=\left\{i:t<n_{i}\leq t^{2}\right\},T_{2}=\left\{i:t^{2}<n_{i}\right\}$ and $T=T_{1}\cup T_{2}$. Define $U_{T_{0}}=\bigotimes_{i\in T_{0}}\operatorname{CZ}_{S_{i}}$ and $U_{T_{1}},U_{T_{2}},U_{T}$ similarly. Without loss of generality, assume $T=[k]$.
The key idea is as follows:
  • For gates in $T_{0}$, we directly use a lightcone argument.
  • For gates in $T_{1}$, we perform a low-degree approximation and use the lightcone argument.
  • For gates in $T_{2}$, we perform a low-degree approximation and use a degree argument.
For each $i$, let $\widetilde{\operatorname{CZ}}_{S_{i}}$ be the low-degree approximation of $\operatorname{CZ}_{S_{i}}$ derived from Corollary 3.10. Let $\tilde{U}_{T_{0}}=\bigotimes_{i\in T_{0}}\operatorname{CZ}_{S_{i}}^{\uparrow},\tilde{U}_{T}=\bigotimes_{i\in T}\widetilde{\operatorname{CZ}}_{S_{i}}^{\uparrow}$ and $\tilde{U}=\tilde{U}_{S}\otimes\tilde{U}_{T}$. Also, let $A^{\Uparrow}$ be the operator dilation of $A$ with respect to the ensemble $\mathcal{S}=\left\{S_{1},\cdots,S_{m}\right\}$. We prove
$\left\|U^{\uparrow}A^{\Uparrow}(U^{\dagger})^{\uparrow}-\tilde{U}A^{\Uparrow}\tilde{U}^{\dagger}\right\|\leq Cn\cdot 2^{-2^{-9}r}.$ (42)
Moreover, we show that the top-left part of $U^{\uparrow}A^{\Uparrow}(U^{\dagger})^{\uparrow}$ is exactly $UAU^{\dagger}$, and taking $M$ to be the top-left part of $\tilde{U}A^{\Uparrow}\tilde{U}^{\dagger}$ satisfies the requirements Eq. 17 and Eq. 18 and will complete the proof.
To begin with, we prove Eq. 42 with a hybrid argument. Let $U_{(i)}=\tilde{U}_{S}\otimes\bigotimes_{j=1}^{i}\widetilde{\operatorname{CZ}}_{S_{j}}^{\uparrow}\otimes\bigotimes_{j=i+1}^{k}\operatorname{CZ}_{S_{j}}^{\uparrow}$. Now, $U_{(0)}=U^{\uparrow}$ and $U_{(k)}=\tilde{U}$. For $1\leq i\leq k$,
$\left\|U_{(i)}-U_{(i-1)}\right\|$ $=\left\|\tilde{U_{S}}\otimes\bigotimes_{j=1}^{i-1}\widetilde{\operatorname{CZ}}_{S_{j}}^{\uparrow}\otimes\bigotimes_{j=i+1}^{k}\operatorname{CZ}_{S_{j}}^{\uparrow}\otimes\left(\operatorname{CZ}_{S_{i}}^{\uparrow}-\widetilde{\operatorname{CZ}}_{S_{i}}^{\uparrow}\right)\right\|$ (43)
$=\left\|\operatorname{CZ}_{S_{i}}^{\uparrow}-\widetilde{\operatorname{CZ}}_{S_{i}}^{\uparrow}\right\|$ (44)
$\leq C_{1}2^{-2^{-9}r}$ (45)
where $C_{1}$ is a constant and the last inequality holds by Corollary 3.10.
Thus,
$\left\|U_{(0)}-U_{(k)}\right\|$ $\leq\sum_{i=1}^{k}\left\|U_{(i)}-U_{(i-1)}\right\|\leq C_{1}n2^{-2^{-9}r}.$ (46)
Hence, choosing $C=2C_{1}$,
$\;\;\left\|U^{\uparrow}A^{\Uparrow}(U^{\dagger})^{\uparrow}-\tilde{U}A^{\Uparrow}\tilde{U}^{\dagger}\right\|$ (47)
$\leq 2\cdot\left\|U^{\uparrow}-\tilde{U}\right\|\cdot\left\|A^{\Uparrow}\right\|$ (48)
$\leq 2\cdot\left\|U^{\uparrow}-\tilde{U}\right\|\cdot\left\|A\right\|$ (49)
$\leq Cn2^{-2^{-9}r}\cdot\left\|A\right\|.$ (50)
where the second inequality uses Lemma 3.8.
For Pauli matrix $\mathcal{B}_{\sigma}$ and some unitary dilation $\left\{M_{i}^{\uparrow}\right\}$ with sets $\left\{S_{i}\right\}$, if $\sigma_{S_{i}}=0_{S_{i}}$, then $M_{i}^{\uparrow}$ acts on the identity matrix. So it gets canceled out with itself:
$\left(M_{i}^{\uparrow}\otimes\mathds{1}\right)\mathcal{B}_{\sigma}^{\Uparrow}\left(\left(M_{i}^{\dagger}\right)^{\uparrow}\otimes\mathds{1}\right)=\mathcal{B}_{\sigma}^{\Uparrow}.$ (51)
If $\sigma_{S_{i}}\neq 0_{S_{i}}$,
$\;\;\left(M_{i}^{\uparrow}\otimes\mathds{1}\right)\mathcal{B}_{\sigma}^{\Uparrow}\left(\left(M_{i}^{\dagger}\right)^{\uparrow}\otimes\mathds{1}\right)$ (52)
[…]
If $\left\{M_{i}\right\}$ are unitary matrices, for both cases, it holds that $\mathcal{B}_{\sigma}^{\Uparrow}=\left(M_{i}\mathcal{B}_{\sigma}M_{i}^{\dagger}\right)^{\Uparrow}.$ Hence,
$U^{\uparrow}A^{\Uparrow}(U^{\dagger})^{\uparrow}=\left(UAU^{\dagger}\right)^{\Uparrow}.$ (55)
As a result, $UAU^{\dagger}$ is the top-left part of $U^{\uparrow}A^{\Uparrow}(U^{\dagger})^{\uparrow}$. Now, we choose $M$ as the top-left part of $\tilde{U}A^{\Uparrow}\tilde{U}^{\dagger}$.
Fix a Pauli matrix $\mathcal{B}_{\sigma}$ with $|\operatorname{supp}\left(\sigma\right)|=d\leq\ell$ in $A$. When $\mathcal{B}_{\sigma}$ acts with $U_{S}$, it becomes $U_{S}^{\dagger}\mathcal{B}_{\sigma}U_{S}$. Since $\operatorname{CZ}_{S_{i}}\mathcal{B}_{\sigma}\operatorname{CZ}_{S_{i}}^{\dagger}=\mathds{1}$ if $S_{i}\cap\operatorname{supp}\left(\sigma\right)=\emptyset$, $U_{S}^{\dagger}\mathcal{B}_{\sigma}U_{S}$ has non-trivial part with degree at most $dn_{i}\leq dt$.
When $\mathcal{B}_{\sigma}$ acts with $U_{T_{1}}$, similar to $U_{S}$, the sets disjoint with $\operatorname{supp}\left(\sigma\right)$ degrade to $\mathds{1}$ in the sense of dilation. Now, we just consider $\widetilde{\operatorname{CZ}}_{S_{i}}\mathcal{B}_{\sigma}\widetilde{\operatorname{CZ}}_{S_{i}}^{\dagger}$, it has degree at most $2d\sqrt{n_{i}r}\leq 2dt\sqrt{r}$ since one qubit can get at most $\sqrt{n_{i}r}$ degree respectively from left side and right side.
When $\mathcal{B}_{\sigma}$ acts with $U_{T_{2}}$, we just add the $\sqrt{n_{i}r}$ to the degree. Recall that $\sum_{i}n_{i}\leq n$, from here, one can get degree at most
$\sum\sqrt{n_{i}r}=\sum n_{i}\sqrt{\frac{r}{n_{i}}}\leq n\sqrt{\frac{r}{t^{2}}}=\frac{n}{t}\sqrt{r}.$ (56)
Combine the three cases together, the degree of $M$ is at most
$2\ell t\sqrt{r}+2\cdot\frac{n}{t}\sqrt{r}=2\sqrt{r}\left(\ell t+n/t\right)\leq 4n^{1/2}\ell^{1/2}r^{1/2}.$ (57)
The last inequality holds due to $t=\sqrt{n/\ell}$.
The spectral norm of the top-left part is bound by the spectral norm of the whole matrix, thus,
$\left\|UAU^{\dagger}-M\right\|\leq Cn\cdot 2^{-2^{-9r}}.$ (58)
Proof of Eq. 37.
Observe that the condition $\left\|Q_{i}-V_{i}S_{i}V_{i}^{\dagger}\right\|\leq\varepsilon_{V}<1/n$ contains that $\left\|Q_{n}\right\|\leq 1+\varepsilon_{V}$. Let $Q_{\leq i}=\prod_{j=1}^{i}Q_{j},V_{\leq i}=\prod_{j=i}^{i}V_{i}\cdot S_{i}\cdot V_{i}^{\dagger}$. With a straightforward calculation,
$\;\;\left\|Q-U\otimes U^{\dagger}\right\|$
$=\left\|S\cdot\prod_{i=1}^{n}Q_{i}-S\cdot\prod_{i=1}^{n}V_{i}\cdot S_{i}\cdot V_{i}^{\dagger}\right\|$
$=\left\|\prod_{i=1}^{n}Q_{i}-\prod_{i=1}^{n}V_{i}\cdot S_{i}\cdot V_{i}^{\dagger}\right\|$
$=\left\|Q_{<n}Q_{n}-V_{<n}Q_{n}+V_{<n}Q_{n}-V_{<n}\left(V_{n}S_{n}V_{n}^{\dagger}\right)\right\|$
$\leq\left\|Q_{n}\right\|\cdot\left\|Q_{<n}-V_{<n}\right\|+\left\|Q_{n}-\left(V_{n}S_{n}V_{n}^{\dagger}\right)\right\|\cdot\left\|V_{<n}\right\|$
$\leq(1+\varepsilon_{V})\cdot\left\|Q_{<n}-V_{<n}\right\|+\varepsilon_{V}$
$\leq(1+\varepsilon_{V})^{n}-1\leq 3n\varepsilon_{V}$
where the first equality comes from Lemma 6.3 and the last inequality comes from the fact that $(1+x)^{n}-1\leq 3nx$ for $x\in(0,1/n)$. ∎